### v1

# Joint Scheduling of DER under Demand Charges: Structure and Approximation

Abstract
We study the joint scheduling of behind-the-meter distributed energy resources (DERs), including flexible loads, renewable generation, and battery energy storage systems, under net energy metering frameworks with demand charges. The problem is formulated as a stochastic dynamic program aimed at maximizing expected operational surplus while accounting for renewable generation uncertainty. We analytically characterize the structure of the optimal control policy and show that it admits a threshold-based form. However, due to the strong temporal coupling of the storage and demand charge constraints, the number of conditional branches in the policy scales combinatorially with the scheduling horizon, as it requires a look-ahead over future states. To overcome the high computational complexity in the general formulation, an efficient approximation algorithm is proposed, which searches for the peak demand under a mildly relaxed problem. We show that the algorithm scales linearly with the scheduling horizon. Extensive simulations using two open-source datasets validate the proposed algorithm and compare its performance against different DER control strategies, including a reinforcement learning-based one. Under varying storage and tariff parameters, the results show that the proposed algorithm outperforms various benchmarks in achieving a relatively small solution gap compared to the theoretical upper bound.
Index Terms:
Battery storage systems, demand charges, distributed energy resources, dynamic programming, energy management, flexible demands, net metering, peak searching.
I Background and Literature
[…]
The transition to a clean, resilient, and sustainable energy system hinges on the optimal coordination of behind-the-meter (BTM) DERs, including distributed generation (DG), flexible demands, and battery energy storage system (BESS) [2]. The operation and coordination of these resources are driven by both economic incentives and technical constraints. Under net energy metering (NEM), prosumers can bi-directionally transact energy and money with the distribution utility (DU), who levy a buy (import) rate or a sell (export) rate based on the net-consumption [3].
With the increasing penetration of new and large electric loads such as electric vehicles, electric ovens, and heat pumps, DUs are increasingly supplementing volumetric charges [^1] Charges based on the total net energy consumption. with demand charges. These are defined as \$/kW fees imposed on the customer’s peak demand recorded over a billing cycle, typically measured in 15-minute or hourly intervals, and are introduced to encourage smoother load profiles and reduce the operational costs of grid maintenance [4]. In the parlance of DU tariff design, this mechanism is referred to as a non-coincident peak demand charge, which levies charges based on each customer’s individual peak demand over a billing period, regardless of when it occurs. This contrasts with a coincident peak demand charge, which charges customers based on their demand contribution during the system-wide peak. Non-coincident demand charges are more commonly employed by DUs due to their administrative simplicity and alignment with cost recovery objectives for distribution infrastructure [5].
Demand charges are a common component of commercial and industrial electricity tariffs and are increasingly being extended to residential customers in some jurisdictions. [^2] Demand charges differ from critical peak pricing, a form of time-varying pricing aimed at incentivizing load shedding during critical periods [6].
Utilities such as Pacific Gas and Electric (PG&E), Consolidated Edison (Con Edison), and Florida Power & Light (FPL) apply demand charges that vary based on customer class, time-of-use periods, and peak-coincident intervals [7, 8, 9]. These charges can constitute a substantial portion of monthly electricity bills, thereby incentivizing customers to strategically manage their load profiles. In response, utilities are adopting demand-based pricing not only to better reflect the cost of capacity provision but also to encourage peak load reduction and grid stability. From a system operator’s perspective, demand charges serve as a proxy for managing distribution-level constraints and capacity costs. Consequently, optimizing DER operations under such pricing schemes is essential for both economic efficiency and system reliability.
With the growing adoption of demand charges, several studies have highlighted the role of BTM DER in helping customers mitigate them [10, 4, 11].
This work focuses on the joint scheduling of BTM DER, including flexible demand, renewable DG, and BESS, under NEM frameworks with differentiated buy and sell rates and demand charges (Fig. 1). A comparative overview of related studies in terms of demand flexibility, storage integration, NEM capabilities, and demand charges is provided in Table I. To the best of our knowledge, this work is the first to co-optimize flexible demand and storage operations under the combined presence of NEM and demand charges.
I-A Related Work
Previous research has explored various angles of home energy management systems (HEMS), particularly in optimizing prosumer energy costs and integrating renewable energy sources. Studies such as [12, 13] utilized dynamic programming to minimize prosumer costs by scheduling BESS operations under inelastic demand, focusing on storage optimization without considering demand flexibility or NEM. Similarly, [14, 15] investigated the co-optimization of demand flexibility and BESS, addressing flexible demands and storage but omitting NEM’s bidirectional energy transactions, which complicates the decision problem.
The integration of NEM, which allows prosumers to buy and sell energy, has been addressed in works like [16, 17, 18, 19]. For instance, [16] optimized storage scheduling under NEM, while [17, 18, 19] jointly optimized storage and flexible demands. However, these studies often neglect demand charges, which significantly complicate the scheduling problem due to their impact on peak load costs.
Recent reinforcement learning (RL) advancements have further enriched the field. Lee et al. [20] introduced a novel approach to multi-residential energy scheduling using federated reinforcement learning, incorporating time-of-use and demand charge tariffs alongside energy storage and renewable sources. In parallel, the work in [21] developed a decision framework for demand-side management and NEM, aimed at peak load reduction and enhanced renewable energy utilization. Their work underscores the importance of managing flexible demands within the context of NEM.
In contrast, while [4, 11] incorporated demand charges into their optimization models, addressing peak power consumption penalties, the assumed free disposal of surplus renewables eliminates NEM’s bidirectional energy transactions, which is a critical limitation given the growing real-world prevalence of such interactions. To provide a broader context, Zainab et al. [22] conducted a comprehensive review of HEMS, encompassing demand response programs, smart technologies, and intelligent controllers.
Despite these advancements, a gap remains in the literature for HEMS that co-optimize flexible demands and storage decisions under NEM and demand charges. Although [4, 11] addressed demand charges and [21] emphasized NEM and demand management, few studies combine all four elements. This work addresses this gap by proposing a HEMS that maximizes prosumer surplus under NEM, incorporating demand flexibility, storage scheduling, and demand charges, as summarized in Table I.
TABLE I: Sample of Related Work on Home Energy Management
Related Work Flexible Demands Storage NEM (Buy/Sell) Demand Charges
[12, 13]
[14, 15]
[16]
[4]
[11]
[17, 18, 19]
[20]
[21]
This Work $\boldsymbol{\checkmark}$ $\boldsymbol{\checkmark}$ $\boldsymbol{\checkmark}$ $\boldsymbol{\checkmark}$
I-B Summary of Results and Contributions
This paper advances research on optimal scheduling of BTM DERs with four key contributions:
  • We develop a stochastic dynamic programming framework that jointly optimizes flexible demand and BESS scheduling under NEM and demand charge constraints, maximizing prosumer operational surplus.
  • We show that the optimal policy maintains a threshold structure consistent with previous work [4] while demonstrating how the integration of multiple factors, such as demand flexibility, battery storage, and demand charges under bi-directional power flow leads to prohibitive computational costs of the optimal policy.
  • To overcome the prohibitive computational costs of the optimal solution, we propose an efficient algorithm that optimally searches for the prosumer’s peak demand under a mildly relaxed problem by leveraging key derived properties from the dynamic programming formulation.
  • Through extensive simulations against baseline approaches and a theoretical maximum, we demonstrate how a computationally efficient algorithm outperforms the benchmarks, including a proximal policy optimization-based reinforcement learning algorithm.
The works in [4, 18] are most closely related to ours, as both adopt a stochastic dynamic programming framework for the prosumer decision problem. Our approach departs from [4] in two key aspects: (i) we model demand as flexible and jointly optimize it with battery storage, and (ii) we allow for bi-directional power and monetary flows between the prosumer and the DU under asymmetric import and export prices. Compared to [18], our model incorporates demand charges in addition to the volumetric charge, introducing stronger temporal coupling across decision stages. Furthermore, unlike both [4, 18], our work proposes an efficient approximation algorithm that achieves near-optimal performance, as validated by simulations against a theoretical upper bound.
I-C Mathematical Notations
[…]
TABLE II: Major variables and parameters (alphabetically ordered)
Symbol Description
$B$ Battery storage capacity.
$c_{t}$ Prosumer’s peak demand before time $t$.
$\bm{d}_{t}$ Vector of consumption bundle at time $t$.
$d_{t}^{k}$ Consumption of device $k$ at time $t$.
$\overline{\bm{d}}_{t}$ Consumption bundle’s upper limit.
$e_{t}$ Battery storage output at time $t$
$\underline{e},\overline{e}$ Battery storage discharging and charging limits.
$g_{t}$ Renewable generation at time $t$.
$\gamma$ Salvage value rate.
$t$ Index of time.
$k$ Index of consumption devices.
$P_{t}(\cdot)$ Payment function at time $t$.
$p$ Peak demand price.
$p^{+},p^{-}$ NEM X buy (retail) and sell (export) rates.
$r_{t}(\cdot)$ Reward function at time $t$.
$\rho$ Battery storage discharging efficiency.
$s_{t}$ Storage state of charge at time $t$.
$\tau$ Battery storage charging efficiency.
$u_{t}$ System control action at time $t$.
$U_{t}(\cdot)$ Utility of consumption at time $t$.
$v_{t}$ Net consumption at time $t$ excluding renewables.
$V^{\pi},V^{\ast}$ Value function under policy $\pi$ and optimal value function.
$x_{t}$ System state.
$z_{t}$ Net consumption at time $t$.
II Problem Formulation
Figure 1: DER setup with flexible demands $d\in\mathbb{R}_{+}$, renewable DG $g\in\mathbb{R}_{+}$, storage operation $e\in\mathbb{R}$, and net-consumption $z\in\mathbb{R}$. The net energy metering tariff with demand charges is also demonstrated with cost associated with amount of net consuming, net producing and the peak load over the period of day where peak load is recorded.
We study a HEMS for prosumers that co-optimizes flexible demand, storage operation, and renewable distributed generation under a NEM tariff with distinct buy/sell rates and peak demand charges. The optimization objective maximizes the operational surplus defined as consumption utility minus electricity costs, where demand charges are levied on the maximum consumption drawn from either load, battery charging, or both. As shown in Fig. 1, the energy storage and flexible loads are jointly scheduled to reduce the demand charge by shaving the load curve, and reduce the volumetric charge by arbitraging NEM rates, while still satisfying prosumer’s consumption.
II-A Prosumer Resources
We consider the operation over a discrete and finite time horizon $t=0,1,\ldots,T$. As shown in Fig.1, the prosumer has BTM renewable DG, denoted by $g_{t}$. The renewable DG is modeled as an exogenous (positive) Markov random process.
The flexible demands are denoted by the vector
$\bm{d}_{t}=(d_{t}^{1},\ldots,d_{t}^{K})\in\mathcal{D}\coloneqq\{\bm{d}:\bm{0}\preceq\bm{d}\preceq\overline{\bm{d}}\}\subseteq\mathbb{R}^{K}_{+},$ (1)
where $K$ is the total number of prosumer devices. Here $\overline{\bm{d}}$ is the consumption bundle’s upper limit.
[…]
The prosumer also owns a BTM battery energy storage, which has output $e_{t}\in[-\underline{e},\overline{e}]$, where $\underline{e}$ and $\overline{e}$ are discharging and charging limits. The output is decomposed into charging $[e_{t}]^{+}=\max\{e_{t},0\}$ and discharging $[e_{t}]^{-}=\max\{-e_{t},0\}$. With charge/discharge efficiencies $\tau,\rho\in(0,1]$, the state of charge (SoC) $s_{t}\in[0,B]$ evolves as:
$s_{t+1}=s_{t}+\tau[e_{t}]^{+}-[e_{t}]^{-}/\rho,\quad t=0,\ldots,T-1,$ (2)
where $B$ denotes battery storage capacity.
The net energy consumption of the prosumer at time $t$ is
$z_{t}:=\bm{1}^{\top}\bm{d}_{t}+e_{t}-g_{t}.$ (3)
Therefore, the prosumer imports power from the grid if $z_{t}>0$ and exports power to the grid if $z_{t}<0$.
II-B System Dynamics
The system state at time $t$, $x_{t}$ is described by $x_{t}=[s_{t},g_{t},c_{t}]$, where $c_{t}$ is the prosumer’s peak demand before time $t$. We allow the system to be controlled by battery charging/discharging and consumption adjustments, i.e., the control action is $u_{t}\coloneqq(e_{t},\bm{d}_{t})$. The peak demand in the state evolves as
$c_{t+1}=\max_{0\leq i\leq t}z_{i}=\max\{z_{t},c_{t}\},\quad t=0,1,\ldots,T-1,$ (4)
and the battery SoC evolves as per (2).
II-C NEM X Tariff with Demand Charges
Under NEM, when the prosumer’s gross loads (from consumption and storage charging) are higher than its generation (from renewable DG), the prosumer pays at the retail rate. Otherwise, the prosumer gets compensated at the export rate.
Let $(p_{t}^{+}\geq p_{t}^{-},p)\geq 0$ be the retail, export, and peak demand prices, respectively. We equivalently break down the peak demand charge at the end of time $T$ to each time step $t$ since $pc_{T}=p\sum_{t=0}^{T-1}[z_{t}-c_{t}]^{+}$. Thus, we can write the payment (cost) under NEM X in [3] for each time $t$ as
$P_{t}(z_{t},c_{t})=p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}+p[z_{t}-c_{t}]^{+},$ (5)
where $P_{t}(\cdot)>0$ ($P_{t}(\cdot)<0$) indicates that the prosumer pays (gets compensated) to (by) the DU.
II-D Prosumer Utility Function
We employ utility functions to formally represent prosumer preferences over consumption bundles, following the standard framework in microeconomic theory [23]. At each time step $t$, the prosumer’s utility of consuming $\bm{d}_{t}$ is given by an additive utility function
$U_{t}(\bm{d}_{t})=\sum_{i=1}^{K}U_{t}(d_{t}^{i}),\quad t=0,1,\ldots,T-1$ (6)
that is assumed to be concave and continuously differentiable. Without loss of generality, we assume $\nabla U_{t}(\bm{d}_{t})\geq 0$, $\bm{d}\in\mathcal{D}$.
II-E DER Co-Optimization Problem
To optimally schedule its DER, the prosumer solves a dynamic program that maximizes its reward function by explicitly accounting for the temporal coupling of storage operations and an NEM tariff with bidirectional energy pricing and peak demand charges.
The reward function is defined as the difference between the utility of consumption (6) and payment under NEM and demand charges (5). Therefore, the reward function is
$r_{t}(x_{t},u_{t})=r_{t}((s_{t},g_{t},c_{t}),(e_{t},\mathbf{d}_{t}))=U_{t}(\bm{d}_{t})-P_{t}(z_{t},c_{t}),\quad t=0,\ldots,T-1.$ (7)
To encourage the preservation of energy in the battery, we define a linear reward to the remaining energy at the terminal stage $T$, i.e.
$r_{T}(x_{T})=r_{T}(s_{T},g_{T},c_{T})=\gamma s_{T},$ (8)
where $\gamma>0$ is the value assigned to the remaining charge in the storage at the end of the horizon.
The policy $\pi=(\pi_{0},\ldots,\pi_{T-1})$ is a sequence of mapping from state space to action space such that $\pi_{t}(x_{t})=u_{t}$. Given initial state $x_{0}=(s,g,0)$, the storage-consumption co-optimization is then defined as:
$\max_{\pi}\quad$ $\mathbb{E}\left[\sum_{t=0}^{T}r_{t}(x_{t},u_{t})\right]$ (9a)
Subject to for $\quad\text{for all}\ t=0,1,\ldots,T-1$
$x_{0}=(s,g,0)$ (9b)
$u_{t}=(e_{t},\bm{d}_{t})=\pi_{t}(x_{t})$ (9c)
$e_{t}\in[-\underline{e},\overline{e}]$ (9d)
$s_{t+1}=s_{t}+\tau[e_{t}]^{+}-[e_{t}]^{-}/\rho$ (9e)
$0\leq s_{t}\leq B$ (9f)
$g_{t+1}\sim F_{\cdot|g_{t}}$ (9g)
$c_{t+1}=\max\{\bm{1}^{\top}\bm{d}_{t}+e_{t}-g_{t},c_{t}\}$ (9h)
$\bm{0}\preceq\bm{d}_{t}\preceq\bm{\overline{d}}.$ (9i)
Constraints (9e) and (9h) induce intertemporal coupling, precluding a myopic (stage-wise optimal) policy. Specifically, relaxing either constraint in isolation is insufficient to eliminate the temporal dependencies, and the resulting optimal policy remains non-myopic (see Thm. 4 in the Appendix). However, when both constraints are simultaneously relaxed, the problem admits a myopic optimal policy characterized by decoupled per-period decisions (see Lem. 4).
The value function $V_{t}^{\pi}(\cdot)$ is defined as the expected total reward received under the policy $\pi$, i.e., $V_{t}^{\pi}(x)=\mathbb{E}[\sum_{t=0}^{T}r_{t}(x_{t},\pi(x_{t}))|x_{t}=x]$. The optimal value function is defined as $V_{t}^{*}(x)=\max_{\pi}V_{t}^{\pi}(x)$.
III Dynamic Program Analytical Results
[…]
The following lemma establishes a conclusion regarding the concavity of the optimal value function.
Lemma 1.
The optimal value function $V^{*}_{t}(s_{t},g_{t},c_{t})$ is concave in $(s_{t},c_{t})$ for every time step $t\leq T$.
Proof.
We follow the proof technique for Lemma 1 in [4].
[…]
First, by the concavity of $U_{t}$:By definition, $U_{t}$ is concave:
$(U_{t}(\bm{d}_{t})+U_{t}(\bm{d}_{t}^{\prime}))/2\leq U_{t}((\bm{d}_{t}+\bm{d}_{t}^{\prime})/{2})=U_{t}(\tilde{\bm{d}}_{t}).$ (10)
[…]
The cost $P_{t}(z_{t},c_{t})$ consists of two components: the energy cost $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ and the demand charging $p[z_{t}-c_{t}]^{+}$. Since $p_{t}^{+}\geq p_{t}^{-}\geq 0$, the energy cost is convex.We separate the $P_{t}(z_{t},c_{t})$ into energy cost $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ and the demand charging cost $p[z_{t}-c_{t}]^{+}$. By assumption, $p_{t}^{+}\geq p_{t}^{-}\geq 0$, and $p_{t}^{+}[0]^{+}=p_{t}^{-}[0]^{-}=0$, the left derivative at 00 is smaller than the right derivative at 00. So, the derivative of $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ is non-decreasing and hence $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ is convex. Then,
$-\sum_{t=\tau}^{T}(p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}+p_{t}^{+}[z_{t}^{\prime}]^{+}-p_{t}^{-}[z_{t}^{\prime}]^{-})/2$ $\leq-\sum_{t=\tau}^{T}p_{t}^{+}[(z_{t}+z_{t}^{\prime})/2]^{+}-p_{t}^{-}[(z_{t}+z_{t}^{\prime})/2]^{-}$ (11)
$=-\sum_{t=\tau}^{T}p_{t}^{+}[\tilde{z}_{t}]^{+}-p_{t}^{-}[\tilde{z}_{t}]^{-},$
and the demand charging cost becomes
$-\sum_{t=\tau}^{T}(p[z_{t}-c_{t}]^{+}+p[z_{t}^{\prime}-c_{t}^{\prime}]^{+})/2=-p(\max_{\tau\leq t\leq T}\{z_{t}\}-c_{\tau}+\max_{\tau\leq t\leq T}\{z_{t}^{\prime}\}-c_{\tau}^{\prime})/2$ $\leq-p(\max_{\tau\leq t\leq T}\{(z_{t}+z_{t}^{\prime})/2\}-(c_{\tau}+c_{\tau}^{\prime})/2)$ (12)
$=-p(\max_{\tau\leq t\leq T}\{\tilde{z}_{t}\}-\tilde{c}_{\tau})=\sum_{t=\tau}^{T}p[\tilde{z}_{t}-\tilde{c}_{t}]^{+}.$
Combining (10)-(12),
$V^{*}_{\tau}(\tilde{x}_{\tau})\geq V_{\tau}^{\tilde{\pi}}(\tilde{x}_{\tau})\geq(V_{\tau}^{*}(x_{\tau})+V_{\tau}^{*}(x_{\tau}^{\prime}))/2,$ (13)
which proves the concavity of the optimal value function. ∎
[…]
It turns out that the optimal value function is also non-decreasing in the system state, which is formally stated in the next lemma.
Lemma 2.
[…]
Proof.
(1) $s_{t}$: Consider two SoCs $s_{t}$, $s_{t}^{\prime}$ such that $s_{t}=s_{t}^{\prime}+\epsilon$, $0<\epsilon\ll\underline{e}$. Let $\{u_{\tau}^{\prime}=(e_{\tau}^{\prime},\bm{d}_{\tau}^{\prime})\}_{\tau=t}^{T}$ be the optimal action sequence for SoC $s_{t}^{\prime}$. If there exists $t\leq\tau_{0}<T$, $e_{\tau_{0}}^{\prime}\geq-\underline{e}+\epsilon$, action sequence $\{u_{\tau}=(e_{\tau}^{\prime}-\epsilon\cdot\mathbf{1}[\tau=\tau_{0}],\bm{d}_{\tau}^{\prime})\}_{\tau=t}^{T}$ is feasible for SoC $s_{t}$. Hence $V^{*}_{t}(s_{t},g_{t},c_{t})-V^{*}_{t}(s_{t}^{\prime},g_{t},c_{t})\geq P_{\tau_{0}}(z_{\tau_{0}}^{\prime},c_{\tau_{0}}^{\prime})-P_{\tau_{0}}(z_{\tau_{0}}^{\prime}-\epsilon,c_{\tau_{0}}^{\prime})\geq 0$. Otherwise, $e_{\tau}^{\prime}<-\underline{e}+\epsilon<0$, $t\leq\tau\leq T$, so action sequence $\{u_{\tau}^{\prime}\}_{\tau=t}^{T}$ is feasible for SoC $s_{t}$, $V^{*}_{t}(s_{t},g_{t},c_{t})-V^{*}_{t}(s_{t}^{\prime},g_{t},c_{t})\geq\gamma(s_{T}-s_{T}^{\prime})=\gamma\epsilon>0$.
(2) $g_{t}$: Consider two generations $g_{t}=g_{t}^{\prime}+\epsilon$, $\epsilon>0$. Let $\pi^{\prime}=\{u_{\tau}^{\prime}=(e_{\tau}^{\prime},\bm{d}_{\tau}^{\prime})\}_{\tau=t}^{T}$ be the optimal action sequence for generation $g_{t}^{\prime}$. Apply the action sequence to generation $g_{t}$, $z_{t+1}=e_{t}^{\prime}+\bm{1}^{\top}\bm{d}_{t}^{\prime}-g_{t}=z_{t+1}^{\prime}-\epsilon$, $V_{t}^{*}(s_{t},g_{t},c_{t})-V_{t}^{*}(s_{t},g_{t}^{\prime},c_{t})\geq P_{t}(z_{t}^{\prime},c_{t})-P_{t}(z_{t}^{\prime}-\epsilon,c_{t})\geq 0$.
(3) $c_{t}$: From (5), $P_{t}(z_{t},c_{t})$ is non-increasing in $c_{t}$, hence $V_{t}^{*}(s_{t},g_{t},c_{t})$ is non-decreasing in $c_{t}$. ∎
[…]
Lastly, we derive a property that links the optimal Q-function, defined as $Q_{t}^{*}(x_{t},u_{t})=r(x_{t},u_{t})+\mathbb{E}[V_{t+1}^{*}(x_{t+1})]$, to the system’s control action $u_{t}$.
Lemma 3.
The optimal Q-function $Q_{t}^{*}(x_{t},u_{t})$ is concave in $u_{t}$.
Proof.
For $u_{\tau}\neq u_{\tau}^{\prime}$, let $\{(x_{t},u_{t})\}_{t=\tau}^{T}$ be the trajectory of $Q_{\tau}^{*}(x_{\tau},u_{\tau})$ and $\{(x_{t}^{\prime},u_{t}^{\prime})\}_{t=\tau}^{T}$ be the trajectory of $Q_{\tau}^{*}(x_{\tau},u_{\tau}^{\prime})$. By (10) - (12), applying actions $\{(u_{t}+u_{t}^{\prime})/2\}_{t=\tau}^{T}$ to $x_{\tau}$ gives:
$Q_{\tau}^{*}(x_{\tau},u_{\tau})+Q_{\tau}^{*}(x_{\tau},u_{\tau}^{\prime})\leq$ $\mathbb{E}\left[\sum_{t=\tau}^{T}r_{t}(x_{t},(u_{t}+u_{t}^{\prime})/2)|x_{\tau}\right]$
$\leq$ $Q_{\tau}^{*}(x_{\tau},(u_{\tau}+u_{\tau}^{\prime})/2).$
Therefore, $Q_{t}^{*}(x_{t},u_{t})$ is concave in $u_{t}$. ∎
Building on the results of Lem. 1, Lem. 2, and Lem. 3, we establish Thm. 1, which characterizes the structure of the optimal co-optimization policy under demand charges.
Theorem 1.
For each step $t=0,1,\ldots,T-1$, the optimal policy is a threshold policy.
Proof.
[…]
$u_{t}^{*}=\arg\max_{u\in\mathcal{F}_{t}}\{r_{t}(x_{t},u)+Q_{t}^{*}(x_{t},u)\}$ (14)
[…]
must exist in the interior of the feasible set $\mathcal{F}_{t}$ or at its boundary. The optimal action can be obtained by solving concave optimizations with and without boundary constraints, which results in a threshold policy. So, we can get the optimal solution by solving $(K+1)2^{K}$ constrained concave optimization and one unconstrained concave optimization, which results in a threshold solution of $(K+1)2^{K}+1$ cases. ∎
Although Thm. 1 reveals a structured policy, direct dynamic programming implementation remains impractical. First, optimization necessitates gradients of $r_{t}$ and $Q_{t}$, which are computationally prohibitive for complex utility functions. Also, continuous-space Q-function tracking requires costly discretization, as validated in simpler settings [4]. To this end, we develop an efficient approximation algorithm to overcome these computational barriers.
IV Efficient Approximation Algorithm
Since computing the exact optimal solution for problem (9) is computationally intractable, we propose an efficient approximation algorithm. The large storage peak searching (LSPS) algorithm efficiently searches for the optimal peak demand by solving a relaxed version of the original problem. The resulting solution is then projected onto the feasible region $\mathcal{F}_{t}$, producing near-optimal and computationally tractable approximations. The original formulation prevents myopic solutions through constraints (9e) and (9h). To address this computational challenge, we first relax constraint (9e) by considering a large capacity regime (Sec. IV-A), then develop an efficient method to handle constraint (9h) (Sec. IV-B).
IV-A Relaxation of the DER Co-Optimization Problem
We consider a large capacity regime as a relaxation of the battery capacity constraint (9e). Observing that $e_{t}$ and $\mathbf{d}_{t}$ consistently appear together in the expression $\mathbf{1}^{\top}\mathbf{d}_{t}+e_{t}$, we introduce a helper function $h_{t}(v_{t})$:
$h_{t}(v_{t}):=\max\{U_{t}(\bm{d}_{t})+\gamma e_{t}:\bm{1}^{\top}\bm{d}_{t}+e_{t}=v_{t},\bm{0}\preceq\bm{d}_{t}\preceq\overline{\bm{d}},-\underline{e}\leq e_{t}\leq\overline{e}\},$ (15)
where $v_{t}$ represents the net consumption excluding renewables at time step $t$. The function $h_{t}(v_{t})$ is differentiable, monotonically increasing, and concave for all time steps $t$. For any $v_{t}\in[-\underline{e},\overline{e}+\bm{1}^{\top}\bm{\overline{d}}]$, there exists a valid pair $(e_{t},\mathbf{d}_{t})$ that achieves $h_{t}(v_{t})$ with $\mathbf{1}^{\top}\mathbf{d}_{t}+e_{t}=v_{t}$. This equivalence allows us to optimize over $v_{t}$ after relaxing constraint (9e).
Using the helper function (15), we formulate the relaxed problem as:
$\max_{v_{0},\ldots,v_{T-1}}\quad$ $\sum_{t=0}^{T-1}h_{t}(v_{t})-p_{t}^{+}[v_{t}-g_{t}]^{+}+p_{t}^{-}[v_{t}-g_{t}]^{-}-p[v_{t}-g_{t}-c_{t}]^{+}$
Subject to $-\underline{e}\leq v_{t}\leq\overline{e}+\bm{1}^{\top}\overline{\bm{d}}$ (16a)
$c_{t+1}=\max\{c_{t},v_{t}-g_{t}\},$ (16b)
To eliminate the temporal dependencies introduced by constraint (9h), we consider a fixed peak demand bound $c$ in the large capacity regime (16):
$\max_{v_{0},\ldots,v_{T-1}}\quad$ $\sum_{t=0}^{T-1}h_{t}(v_{t})-p_{t}^{+}[v_{t}-g_{t}]^{+}+p_{t}^{-}[v_{t}-g_{t}]^{-}-pc$ (17a)
Subject to $-\underline{e}\leq v_{t}\leq\overline{e}+\bm{1}^{\top}\overline{\bm{d}},$ (17b)
$v_{t}\leq g_{t}+c.$ (17c)
This reformulation replaces the recursive peak demand update (16b) with a single-step upper-bound constraint (17c), effectively removing temporal dependencies while preserving the optimal policy structure. We show that the optimal policy of (16) is preserved under this relaxation.
Theorem 2.
Suppose $\{v_{t}^{*}\}$ and $c^{*}$ are optimal actions and peak demand for (16), then $\{v_{t}^{*}\}$ is also optimal for (17) with demand bound $c=c^{*}$.
Proof.
Suppose $\{v_{t}^{*}\}$ is the optimal solution for problem (16). Since $c^{*}$ is the optimal peak demand, we have $v_{t}-g_{t}\leq c^{*}=c$ for all $t$. This means $\{v_{t}\}$ is also feasible for problem (17). Since problem (17) searches within a subspace of problem (16), $\{v_{t}^{*}\}$ is also optimal for problem (17). ∎
By Thm. 2, we can compute the optimal policy for (16) by searching for the optimal $c^{*}$ and solving (17).
IV-B Solution under Given Peak Demand
With a fixed peak demand bound $c$, the optimization problem decomposes into independent single-step subproblems that can be solved analytically. Due to the concavity of $h_{t}$, we first derive the unconstrained optimum $v^{\dagger}_{t}(g_{t})$, then project it onto the feasible region $\mathcal{F}_{t}$ defined by the constraints to obtain the constrained optimum $v^{*}_{t}(g_{t})$.
Lemma 4.
For $t=0,\ldots,T-1$, the optimal policy for (17) is given by
$v_{t}^{\ast}(g_{t})=\begin{cases}-\underline{e},\quad\quad\quad\quad\quad\ v^{\dagger}(g_{t})\leq-\underline{e},\\ v_{t}^{\dagger}(g_{t}),\quad\quad\quad\quad-\underline{e}\leq v^{\dagger}(g_{t})\leq\min\{c+g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}\},\\ \min\{c+g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}\},\min\{c+g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}\}\leq v^{\dagger}(g_{t}),\end{cases}$ (18)
where the unconstrained optimum $v_{t}^{\dagger}(g_{t})$ is the solution without constraints (17b) and (17c), given by
$v_{t}^{\dagger}(g_{t})=\begin{cases}(h^{\prime})^{-1}(p_{t}^{-}),&(h^{\prime})^{-1}(p_{t}^{-})<g_{t},\\ g_{t},&(h^{\prime})^{-1}(p_{t}^{+})\leq g_{t}\leq(h^{\prime})^{-1}(p_{t}^{-}),\\ (h^{\prime})^{-1}(p_{t}^{+}),&g_{t}<(h^{\prime})^{-1}(p_{t}^{+}).\end{cases}$ (19)
Proof.
With fixed peak demand, the optimization decomposes into independent subproblems at each time step $t=0,1,\cdots,T-1$:
$\max_{v_{t}}\quad$ $h_{t}(v_{t})-p_{t}^{+}[v_{t}-g_{t}]^{+}+p_{t}^{-}[v_{t}-g_{t}]^{-}$ (20a)
Subject to $v_{t}\in[-\underline{e},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}],\quad v_{t}-g_{t}\leq c.$ (20b)
The unconstrained optimum is
$v_{t}^{\dagger}(g_{t})=\arg\max_{v}~{}~{}h_{t}(v)-p_{t}^{+}[v-g_{t}]^{+}+p_{t}^{-}[v-g_{t}]^{-}=\begin{cases}(h_{t}^{\prime})^{-1}(p_{t}^{-}),&(h_{t}^{\prime})^{-1}(p_{t}^{-})<g_{t},\\ g_{t},&(h_{t}^{\prime})^{-1}(p_{t}^{+})<g_{t}<(h_{t}^{\prime})^{-1}(p_{t}^{-}),\\ (h_{t}^{\prime})^{-1}(p_{t}^{+}),&g_{t}<(h_{t}^{\prime})^{-1}(p_{t}^{+}).\end{cases}$
Projecting onto the feasible region defined by constraint (20b) yields the optimal policy in (18). ∎
The optimal policy has a clear economic interpretation based on the relationship between renewable generation and consumption needs. The unconstrained solution $v_{t}^{\dagger}(g_{t})$ balances the marginal utility of consumption against energy trading costs. When $(h_{t}^{\prime})^{-1}(p_{t}^{-})<g_{t}$, the renewable generation $g_{t}$ exceeds the energy demanded by consuming at the selling price $\pi^{-}$ and charging the battery. In this case, the optimal consumption $v_{t}^{\dagger}(g_{t})=(h_{t}^{\prime})^{-1}(p_{t}^{-})$ is set such that the marginal utility equals the selling price, meaning excess generation beyond this consumption level should be sold to the grid. When $(h_{t}^{\prime})^{-1}(p_{t}^{+})\leq g_{t}\leq(h_{t}^{\prime})^{-1}(p_{t}^{-})$, the renewable generation exactly matches the optimal consumption level, so $v_{t}^{\dagger}(g_{t})=g_{t}$ and no energy trading occurs. When $g_{t}<(h_{t}^{\prime})^{-1}(p_{t}^{+})$, renewable generation is insufficient relative to the consumption level that would justify purchasing energy at price $p_{t}^{+}$. Here, the optimal consumption $v_{t}^{\dagger}(g_{t})=(h_{t}^{\prime})^{-1}(p_{t}^{+})$ sets marginal utility equal to the purchase price, requiring additional energy to be bought from the grid. The final policy $v_{t}^{*}(g_{t})$ then projects this unconstrained optimum onto the feasible region, ensuring that consumption respects both battery operation limits and peak demand constraints.
IV-C Algorithmic Search of the Optimal Peak Demand
We leverage the concavity properties of the objective function to efficiently determine the optimal peak demand $c^{*}$ without solving the complete optimization problem (16). Our approach exploits the structural properties of the relaxed co-optimization problem (17).
Lemma 5.
Let $J(c)$ be the optimal objective function of (17),
$J(c):=\sum_{t=0}^{T-1}h_{t}(v_{t}^{*})-p_{t}^{+}[v_{t}^{*}-g_{t}]^{+}+p_{t}^{-}[v_{t}^{*}-g_{t}]^{-}-pc,$ (21)
where $v_{t}^{*}$ depends on $c$ through the optimization. Then $J(c)$ is concave in $c$.
Proof.
Let $I(c)=\{t:v_{t}^{*}=c+g_{t}\}$ be the set of time indices where the peak demand constraint is active. We can decompose the objective function as
$J(c)=-pc+\sum_{t\in I(c)}h_{t}(c+g_{t})-p_{t}^{+}c+\sum_{t\in[T-1]\setminus I(c)}h_{t}(v_{t}^{*})-p_{t}^{+}[v_{t}^{*}-g_{t}]^{+}+p_{t}^{-}[v_{t}^{*}-g_{t}]^{-}.$
As $c$ increases, the constraint set becomes less restrictive, so $|I(c)|$ is non-increasing. For the terms not in $I(c)$, the optimal $v_{t}^{*}$ is independent of $c$ and determined by the unconstrained optimum. Taking the first and second derivatives,
$J^{\prime}(c)=-p+\sum_{t\in I(c)}h^{\prime}_{t}(c+g_{t})-p_{t}^{+},$
$J^{\prime\prime}(c)=\sum_{t\in I(c)}h^{\prime\prime}_{t}(c+g_{t})\leq 0,$
where the inequality follows from the concavity of $h_{t}$. Therefore, $J(c)$ is concave in $c$. ∎
The concavity of $J(c)$ ensures that the optimal $c^{*}$ is unique and can be found by solving the first-order condition $J^{\prime}(c^{*})=0$. The key insight is that $I(c)$ changes only at discrete values of $c$, allowing us to search for the optimum efficiently. Specifically, define the candidate set
$\mathcal{C}=\{\min\{v_{t}^{\dagger}(g_{t})-g_{t},\bm{1}^{\top}\overline{\bm{d}}+\overline{e}-g_{t}\}\}_{t=0}^{T-1},$
where $v_{t}^{\dagger}(g_{t})$ is the unconstrained optimum from (19). The set $I(c)$ remains constant for $c\in[c_{1},c_{2}]$, where
$c_{1}=\max\{c:c\in\mathcal{C},J^{\prime}(c)<0\},$ (22)
$c_{2}=\min\{c:c\in\mathcal{C},J^{\prime}(c)>0\}.$
The optimal $c^{*}$ lies in this interval and satisfies
$J^{\prime}(c^{*})=-p+\sum_{t\in I(c_{1})}h^{\prime}_{t}(c^{*}+g_{t})-p_{t}^{+}=0.$ (23)
IV-D Large Storage Peak Searching (LSPS) Algorithm
The LSPS algorithm, in Alg. 1, efficiently approximates the problem (9) by combining the relaxation techniques from the previous sections. It leverages the key structural properties established in the previous sections: the decomposition enabled by the large capacity relaxation, the analytical solution for fixed peak demand, and the concavity of the objective function in the peak demand parameter. By searching over the discrete candidate set $\mathcal{C}$, we avoid solving the full dynamic optimization problem and rather solve a linear complexity algorithm, as formalized below, while maintaining a small solution gap compared to the optimal.
Algorithm 1 Large storage peak searching (LSPS) algorithm
1:Candidate set $\mathcal{C}\leftarrow\{\min\{v_{t}^{\dagger}(g_{t})-g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}-g_{t}\}\}_{t=0}^{T-1}$
2:Derivatives $J^{\prime}(\mathcal{C})\leftarrow\{J^{\prime}(c):c\in\mathcal{C}\}$
3:$c_{1}\leftarrow\max\{c:c\in\mathcal{C},J^{\prime}(c)<0\}$
4:Solve $J^{\prime}(c^{*})=0$ using (23) for optimal peak demand $c^{*}$
5:Find optimal actions $v_{t}^{*}$ using (18) with peak demand $c^{*}$
6:Project $v_{t}^{*}$ back to variables $(e_{t}^{*},\mathbf{d}_{t}^{*})$ using (15)
7:Clip $e_{t}^{*}$ to ensure $s_{t+1}\in[0,B]$
Theorem 3.
Alg. 1 has computational complexity $\mathcal{O}(T)$.
Proof.
Under the assumption that $h_{t}^{\prime}$ and $(h_{t}^{\prime})^{-1}$ can be evaluated in constant time, each step of the algorithm requires $\mathcal{O}(T)$ operations. Step 1 computes $T$ candidate values, Step 2 evaluates $T$ derivatives, Step 3 performs a linear search through $\mathcal{O}(T)$ values, Step 4 solves a single equation with less than $T$ terms in $\mathcal{O}(T)$ time, and Steps 5-7 compute $T$ optimal actions with each one constant time. Therefore, the overall complexity is $\mathcal{O}(T)$. ∎
Empirical analysis on the performance of the LSPS algorithm is presented in Sec. V.
V Case Study
Using two rich datasets, and considering a one-day operation horizon $T=24$, we evaluate the DER co-optimization performance of the proposed LSPS algorithm and an RL controller using a Proximal Policy Optimization (PPO) algorithm compared to a theoretical maximum that assumes perfect knowledge of future renewables solving Eq. 9. The performance is measured by the solution gap, which is the difference between the cumulate reward of the algorithms and the theoretical upper bound. Experiments were run on a server with 1 AMD Ryzen Threadripper 3990X 64-Core Processor and 4 Nvidia RTX A4000 GPUs.
V-A Dataset Description
(a) Daily net demand for HouseZero (Jun 2022 – May 2024)
(b) Daily net demand for BDGP building (Jan 2016 – Dec 2017)
Figure 2: Daily net demand distributions for HouseZero and BDGP over the training and testing period.
Two open-source datasets from Harvard HouseZero [24] and Building Data Genome Project (BDGP) [25] were used to evaluate the performance of DER scheduling under different algorithms. The two buildings were chosen with distinct characteristics while both are equipped with solar PV. Renewable generation and demand data, resampled to hourly granularity, are used for training and testing the performance of different control strategies. Fig. 2 depicts the net demand distribution for HouseZero and BDGP buildings. The two buildings differ significantly in their energy pattern: HouseZero is a low-energy, ultra-efficient building with a peak demand below 2 kW and a frequently negative net consumption. In contrast, the BDGP building has substantially higher consumption, with net consumption peaking at over 160 kW.
V-B Simulation Setup
We adopt the concave utility function from [26]: $U_{t}(d_{t})=\alpha_{t}d_{t}-\frac{1}{2}\beta_{t}d_{t}^{2}$, where $\alpha_{t}$ and $\beta_{t}$are time-varying and learned apriori using historical energy prices, demands, and demand elasticity [27]. The elasticity was set to -0.1 [28]. To get more crisp insights, we assumed time-invariant NEM parameters, with a $p^{+}=\$0.12kWh$ retail (import) rate and a $p^{-}=\$0.06/kWh$ sell (export) rate. The daily peak demand charge was set to $p=\$10/kW$. The final salvage value for the remaining energy in the battery was set at $\gamma=\$0.09/kWh$. The baseline battery parameters were selected based on the demand data and size of the building. For HouseZero, $B=5$ kWh and $\underline{e}=\overline{e}=1$ kW were used; while for BDGP building, $B=350$ kWh and $\underline{e}=\overline{e}=50$ kW were used. In both cases, the battery charging/discharging efficiencies were set at $\tau=\rho=0.95$.
V-C Control Strategies
We present five control strategies to schedule DER here:
Backup Mode
The backup mode prioritizes using renewables to charge the battery to maintain a high SoC level, which is usually set by the user. The battery discharges only during grid outages, equivalent to the case without a battery system [4] during normal operations.
Renewable-Adjusted Threshold Policy (RATP)
RATP is a threshold-based battery scheduling policy that adjusts charging and discharging based on the renewable adjusted gross consumption ($\tilde{d}=d-g$). When $\tilde{d}>0$, the battery discharges to reduce grid purchases; when $\tilde{d}<0$, it charges to avoid exporting excess renewables at the low $p^{-}$. When $\tilde{d}=0$ the battery stays idle. Although simple, RATP is highly effective under NEM structures, as it arbitrages the difference between the retail and sell rates by minimizing power imports and exports. However, RATP’s performance significantly degrades under NEM with demand charges, as it is agnostic to strategic net consumption decisions that shape demand peaks.
Mathematically, the storage control under RATP is
$e_{t}(g_{t})=\begin{cases}\max\{-\underline{e},\rho s_{t},-\tilde{d}_{t}\},&\bar{d}_{t}-g_{t}>0\\ \min\{\overline{e},(B-s_{t})/\tau,-\tilde{d}_{t}\}&\bar{d}_{t}-g_{t}\leq 0,\end{cases}$
and $d_{t}=\bar{d}_{t}$ in both cases. We use RATP and Threshold interchangeably.
RL Algorithm
We use a popular model-free RL method of proximal policy optimization (PPO) algorithms to directly learn the optimal policy for controlling the DER system described. PPO uses an actor-network and a critic network with a multilayer perceptron (MLP) structure. For training, each of the actor and critic networks consists of two hidden layers with 256 neurons per layer. The actor-network predicts the optimal action, while the critic network estimates the optimal value function. We train the networks with a learning rate of $1e^{-4}$, batch size of 64, and 16 epochs per update. Other training parameters are kept as default provided by Stable-baselines3 [29].
Large Storage Peak Searching (LSPS) Algorithm
LSPS algorithm, described in Alg. 1, offers an efficient and scalable method to co-optimize DER. Following Alg. 1, we first calculate the optimal peak, then, by solving Eq. 17 with the optimal peak, we get the battery and demand actions at each time step.
Theoretical Upper Bound
The theoretical upper bound assumes perfect knowledge of the renewables for the entire scheduling period and solves the constrained convex program in Eq. 9.
V-D Simulation Results
We compare the performance of the control strategies (a)-(d) above to the theoretical optimal algorithm in (e) by varying the following parameters:
  • Battery capacity ($B$): 5kWh – 50kWh for HouseZero and 350kWh – 950kWh for BDGP building to simulate battery sizes from small to large.
[…]
  • Export rate of electricity ($p^{-}$): \$$\$0/\text{kWh}$\$$\$0.12/\text{kWh}$ from no to high sell back incentive with $p^{-}=p^{+}$.
  • Peak demand price of electricity ($p$): \$$\$0/\text{kW}$\$$\$10/\text{kW}$ from no penalty on a high peak to a reasonably high peak charge commonly practiced in electricity tariffs.
Tab. III summarizes the overall performance comparisons in terms of the solution gap to the optimal solutions. [^3] Table IV in the Appendix details all cases for reference. Overall, the LSPS algorithm performed closest to the optimal solutions over all scenarios with an average gap of 13.77%, followed by RL of 20.21%, threshold algorithm of 41.93%, and backup mode of 43.75%. The LSPS algorithm was closest to the optimal 34 out of 42 scenarios (81%) while RL was closest to optimal in the remaining 8 scenarios (19%). Furthermore, for almost all cases, control strategies/algorithms showed better performance for the smaller building of HouseZero, with LSPS performing only 4.52% below the theoretical optimal. The higher peak demand and energy consumption presented greater challenges for optimal operations of the DER system.
TABLE III: Summary of different control strategies under test cases
Case Building Percentage gaps to optimal
Backup Threshold RL LSPS
Battery capacity HouseZero 47.44% 14.17% 9.33% 1.03%
BDGP 64.42% 63.56% 29.50% 25.37%
Salvage value rate HouseZero 41.06% 52.41% 19.54% 6.06%
BDGP 50.16% 66.56% 29.28% 18.64%
Export rate HouseZero 50.83% 46.21% 22.58% 8.59%
BDGP 62.23% 61.26% 32.76% 31.91%
Peak demand price HouseZero 17.86% 15.85% 8.01% 2.41%
BDGP 16.04% 15.39% 10.70% 8.86%
Overall HouseZero 39.30% 32.16% 14.86% 4.52%
BDGP 48.21% 51.69% 25.56% 21.20%
(a) HouseZero, battery capacity case
(b) HouseZero, salvage value rate case
(c) HouseZero, export rate case
(d) HouseZero, peak charge case
(e) BDGP, battery capacity case
(f) BDGP, salvage value rate case
(g) BDGP, export rate case
(h) BDGP, peak charge case
Figure 3: Comparison of daily surplus gaps to theoretical upper bound for different control strategies with varying parameters using HouseZero (top row) and BDGP (bottom row) test data. The cases include (a,e) battery capacities, (b,f) salvage value rates, (c,g) electricity export rates, and (d,h) peak demand prices.
Illustrated in Fig. 3, for the first group of test cases varying battery capacity sees the LSPS algorithm performed closest to the theoretical upper bound for both HouseZero and BDGP in Fig. 3(a) and Fig. 3(e). For the HouseZero case with low loads and large battery capacities, the peak was completely removed (0kW), leading the LSPS algorithm to perform as well as the theoretical optimal. Large battery capacity also significantly improved the threshold algorithm performance, even though it significantly fell behind RL and LSPS algorithms when the battery size was more limited. On the other hand, RL followed a similar trend as the LSPS algorithm while performing around 10% worse over all cases. RL was still stable and more adaptive to changes in the environment than the threshold and backup algorithms. For the much larger BDGP building, both RL and LSPS algorithms improved overall with the increased battery capacity while backup and threshold failed to show the expected improvement due to their inflexibility in considering the high peak and energy use in this building.
Second, varying salvage value rate again showed the superior performance of the LSPS algorithm for both HouseZero and BDGP in Fig. 3(b) and Fig. 3(f). The threshold algorithm tended to use up the available battery charge, leading to low salvage value at the end of the day and causing a larger gap to optimal with increased salvage value rate. Meanwhile, for extremely high salvage value cases, the backup case got closer to optimal as the strategy kept the battery full at the end of the day. RL performed relatively well and was closer to the LSPS algorithm and optimal for the relatively smaller HouseZero building. Overall, for both HouseZero and BDGP, the performances exhibited a similar trend across the four control strategies/algorithms.
Third, varying export rates highlighted the difference between HouseZero and BDGP buildings illustrated in Fig. 3(c) and Fig. 3(g). For the BDGP building, the higher energy demand meant the building would never be in net producing regime, leading to flat lines for the control strategies/algorithms. With RL, there were slight variations in performance from the stochasticity during training. For HouseZero, all control strategies/algorithms improved with a higher export rate since it generated more revenue for the building to sell back excessive energy production from solar. LSPS and RL algorithms performed relatively well and were stable across different export rates over the two test buildings.
Lastly, varying peak demand charges saw all control strategies/algorithms change in a similar trend shown in Fig. 3(d) and Fig. 3(h). It was observed that for low peak charges, all control strategies/algorithms performed well, highlighting the complexity of peak demand charges added to optimally managing such a DER system. With higher peak demand charges, all control strategies/algorithms diverged from optimal, with LSPS and RL algorithms performing the best across different scenarios. Additionally, both of them showed a smaller solution gap for the smaller building of HouseZero, indicating the challenges for buildings with higher demand and energy consumption.
Overall, both LSPS and RL showed great robustness and performance across different scenarios with varying battery capacities, salvage values, and prices of electricity. Specifically, LSPS optimization outperformed backup and threshold algorithms by 68.47% and 68.9% while RL outperformed them by 52.49% and 48.01% respectively. LSPS and RL algorithms trended in sync in their gap to optimality, demonstrating their ability to capture system dynamics while the LSPS algorithm is superior in around 81% instances. Moreover, LSPS was optimal with a large battery capacity and close to optimal for low export rate and peak demand charge. With 0 or low peak demand charges, all control strategies/algorithms performed close to optimal. In comparison, the threshold algorithm and backup mode were sensitive to changes in battery capacity, salvage rate, peak demand charge, and export rate when the building entered the net-producing regime. For cases where the battery capacity was large or the peak price was low, the threshold algorithm outperformed RL and approaches optimal with deterministic predictions. However, with a high salvage value and demand price, the threshold algorithm’s performance degraded significantly and became even worse than the backup mode. Backup mode varied mostly linearly in its gap to the optimal case. It performed well only in special cases with high salvage value rates and low peak demand cases.
(a) Battery action comparison
(b) Demand action comparison
Figure 4: Typical day battery and demand actions comparison between different control methods for test building HouseZero.
As an example, Fig. 4 showed the action sequence for the baseline configuration for HouseZero over one test day. All strategies followed a similar behavior which discharged the battery at the beginning and end of the day. At the same time, demand was reduced over the evening time when there were no solar generations. Both threshold and RL were more aggressive in battery charging actions when lacking generation predictions. The optimal battery actions indicated a much slower charging behavior over the daytime with precise knowledge of solar radiation, creating the performance gap mentioned above. It was also observed that the LSPS algorithm may generate aggressive demand reduction actions due to the lack of predictions. Such behavior negatively impacted the utility function and the overall objective, causing gaps in the optimal solutions.
VI Conclusion
This work provides the first analytical characterization of the joint optimization of flexible demand and BESS under demand charges and bi-directional power flow. We model the problem using a stochastic dynamic program and show that the optimal policy exhibits a threshold structure. To address computational complexity, we develop an efficient demand peak searching approximation method, based on a relaxed version of the problem, leading to a linear-complexity algorithm that identifies the optimal peak demand.
Through simulations with real data from two rich datasets, we demonstrated the performance of the LSPS algorithm in achieving a small solution gap compared to baseline approaches, highlighting its effectiveness compared to relatively more computationally prohibitive methods such as model-predictive control and RL.
The work presented here has several limitations that open avenues for future research. Firstly, the RL algorithm used in our simulations is relatively simple. Designing a more sophisticated algorithm that reliably converges to the theoretical upper bound remains an important direction for future work. Second, the proposed algorithm assumes full knowledge of the utility’s functional form $U(\cdot)$, which may be difficult to obtain in practice. This limitation can be addressed using machine learning or RL techniques that infer the utility function from observed behavior [30].

### v2

@@
 Abstract
-We study the joint scheduling of behind-the-meter distributed energy resources (DERs), including flexible loads, renewable generation, and battery energy storage systems, under net energy metering frameworks with demand charges.
+We study the joint scheduling of behind-the-meter distributed energy resources (DERs), including flexible loads, renewable generation, and battery energy storage systems, under net energy metering tariffs with demand charges.
 The problem is formulated as a stochastic dynamic program aimed at maximizing expected operational surplus while accounting for renewable generation uncertainty.
-We analytically characterize the structure of the optimal control policy and show that it admits a threshold-based form.
+We analytically characterize the optimal control policy and show that it admits a threshold-based structure.
 However, due to the strong temporal coupling of the storage and demand charge constraints, the number of conditional branches in the policy scales combinatorially with the scheduling horizon, as it requires a look-ahead over future states.
@@
 Extensive simulations using two open-source datasets validate the proposed algorithm and compare its performance against different DER control strategies, including a reinforcement learning-based one.
-Under varying storage and tariff parameters, the results show that the proposed algorithm outperforms various benchmarks in achieving a relatively small solution gap compared to the theoretical upper bound.
-Index Terms:
-Battery storage systems, demand charges, distributed energy resources, dynamic programming, energy management, flexible demands, net metering, peak searching.
+Under varying storage and tariff parameters, the results show that the proposed algorithm outperforms various benchmarks in achieving a relatively small solution gap compared to a theoretical upper bound.
 I Background and Literature
-[…]
 The transition to a clean, resilient, and sustainable energy system hinges on the optimal coordination of behind-the-meter (BTM) DERs, including distributed generation (DG), flexible demands, and battery energy storage system (BESS) [2].
@@
 [^2] Demand charges differ from critical peak pricing, a form of time-varying pricing aimed at incentivizing load shedding during critical periods [6].
-Utilities such as Pacific Gas and Electric (PG&E), Consolidated Edison (Con Edison), and Florida Power & Light (FPL) apply demand charges that vary based on customer class, time-of-use periods, and peak-coincident intervals [7, 8, 9].
+Utilities such as Pacific Gas and Electric (PG&E), Consolidated Edison (Con Edison), Florida Power & Light (FPL), and Arizona Public Service (APS) apply demand charges that vary based on customer class, time-of-use periods, and peak-coincident intervals [7, 8, 9, 10].
 These charges can constitute a substantial portion of monthly electricity bills, thereby incentivizing customers to strategically manage their load profiles.
@@
 However, these studies often neglect demand charges, which significantly complicate the scheduling problem due to their impact on peak load costs.
-Recent reinforcement learning (RL) advancements have further enriched the field.
-Lee et al.
-[20] introduced a novel approach to multi-residential energy scheduling using federated reinforcement learning, incorporating time-of-use and demand charge tariffs alongside energy storage and renewable sources.
-In parallel, the work in [21] developed a decision framework for demand-side management and NEM, aimed at peak load reduction and enhanced renewable energy utilization.
-Their work underscores the importance of managing flexible demands within the context of NEM.
 In contrast, while [4, 12] incorporated demand charges into their optimization models, addressing peak power consumption penalties, the assumed free disposal of surplus renewables eliminates NEM’s bidirectional energy transactions, which is a critical limitation given the growing real-world prevalence of such interactions.
@@
 [21] conducted a comprehensive review of HEMS, encompassing demand response programs, smart technologies, and intelligent controllers.
+Recent progress in reinforcement learning (RL) has further enriched the area by enabling data-driven approaches to jointly optimize DER under complex tariffs.
+For example, Lee et al.
+[22] introduced a federated RL framework for multi-residential energy scheduling that incorporates time-of-use and demand charge tariffs alongside energy storage and renewable sources.
+Similarly, [23] developed a decision framework for demand-side management under NEM, aiming at peak load reduction and improved renewable utilization.
+While RL offers scalability and adaptability to uncertain environments, it generally lacks the interpretability and optimality guarantees of dynamic programming, which can make it difficult to validate solutions or ensure consistent performance across varying operating conditions.
 Despite these advancements, a gap remains in the literature for HEMS that co-optimize flexible demands and storage decisions under NEM and demand charges.
 Although [4, 12] addressed demand charges and [23] emphasized NEM and demand management, few studies combine all four elements.
-This work addresses this gap by proposing a HEMS that maximizes prosumer surplus under NEM, incorporating demand flexibility, storage scheduling, and demand charges, as summarized in Table I.
+This work addresses this gap by proposing a HEMS that maximizes prosumer surplus under NEM, incorporating demand flexibility, storage scheduling, and demand charges, as summarized in Table I, and explained in the next section.
 TABLE I: Sample of Related Work on Home Energy Management
@@
 [23]
-This Work $\boldsymbol{\checkmark}$ $\boldsymbol{\checkmark}$ $\boldsymbol{\checkmark}$ $\boldsymbol{\checkmark}$
+[…]
 I-B Summary of Results and Contributions
@@
   • We develop a stochastic dynamic programming framework that jointly optimizes flexible demand and BESS scheduling under NEM and demand charge constraints, maximizing prosumer operational surplus.
-  • We show that the optimal policy maintains a threshold structure consistent with previous work [4] while demonstrating how the integration of multiple factors, such as demand flexibility, battery storage, and demand charges under bi-directional power flow leads to prohibitive computational costs of the optimal policy.
+  • We show that the optimal policy maintains a threshold structure consistent with previous work [4] while demonstrating how the integration of multiple factors, such as demand flexibility, battery storage, and demand charges under bi-directional power flow, leads to prohibitive computational costs of the optimal policy.
   • To overcome the prohibitive computational costs of the optimal solution, we propose an efficient algorithm that optimally searches for the prosumer’s peak demand under a mildly relaxed problem by leveraging key derived properties from the dynamic programming formulation.
@@
 Our approach departs from [4] in two key aspects: (i) we model demand as flexible and jointly optimize it with battery storage, and (ii) we allow for bi-directional power and monetary flows between the prosumer and the DU under asymmetric import and export prices.
+Symmetric NEM import and export prices, i.e., NEM 1.0, are known to create utility revenue shortfalls [24] potentially leading to death spirals [25] and cross-subsidies [26].
+Consequently, several states, including California, Arizona, Illinois, Georgia, and Utah, have adopted successor NEM tariffs featuring asymmetric import and export prices [27].
 Compared to [19], our model incorporates demand charges in addition to the volumetric charge, introducing stronger temporal coupling across decision stages.
-Furthermore, unlike both [4, 18], our work proposes an efficient approximation algorithm that achieves near-optimal performance, as validated by simulations against a theoretical upper bound.
-I-C Mathematical Notations
-[…]
+From an economic perspective, demand charges improve the cost reflectivity of the tariff and reduce the over-reliance on volumetric charges to recover bundled cost components [24].
+Lastly, unlike both [4, 19], our work proposes an efficient approximation algorithm that outperforms various benchmarks in achieving a relatively small solution gap compared to a theoretical upper bound.
+I-C Mathematical Notations and Paper Organization
+Table II lists the major symbols used.
+Throughout the paper, boldface letters indicate column vectors as in $\bm{x}=(x_{1},\ldots,x_{n})$ with $\bm{x}^{\top}$ as its transpose.
+In particular, $\bm{1}$ is a column vector of all ones.
+For a multivariate differentiable function $f$ of $\bm{x}$, we use interchangeably $f(\bm{x})$ and $f(x_{1},\ldots,x_{n})$, and we use $f^{\prime}$ and $f^{\prime\prime}$ to denote its first-order and second-order derivative (if it exists), respectively.
+We use $f^{-1}$ to denote its inverse function (if it exists).
+For vectors $\bm{x},\bm{y}$, $\bm{x}\preceq\bm{y}$ is the element-wise inequality $x_{i}\leq y_{i}$ for all $i$, and $\bm{x}]^{+},[\bm{x}]^{-}$ are the element-wise positive and negative parts of vector $\bm{x}$, i.e., $x_{i}]^{+}:=\max\{0,x_{i}\}$, $x_{i}]^{-}:=-\min\{0,x_{i}\}$ for all $i$, and $\bm{x}=[\bm{x}]^{+}-[\bm{x}]^{-}$.
+We denote by $\mathbb{R}_{+}:=\{x\in\mathbb{R}:x\geq 0\}$ the set of non-negative real numbers.
+The rest of the paper is organized as follows.
+§II, provides a mathematical formulation of DERs, tariff under demand charge, system dynamics, and the joint scheduling decision problem.
+§III presents some analytical results on the dynamic programming problem, and §IV proposes an efficient approximation algorithm that circumvents the exponential complexity of solving the dynamic program.
+A detailed case study using real data from two buildings is presented in §V, followed by concluding remarks and future directions in §VI.
 TABLE II: Major variables and parameters (alphabetically ordered)
@@
 $d_{t}^{k}$ Consumption of device $k$ at time $t$.
-$\overline{\bm{d}}_{t}$ Consumption bundle’s upper limit.
+$\overline{\bm{d}}$ Consumption bundle’s upper limit.
 $e_{t}$ Battery storage output at time $t$
@@
 $\gamma$ Salvage value rate.
-$t$ Index of time.
 $k$ Index of consumption devices.
@@
 $s_{t}$ Storage state of charge at time $t$.
+$t$ Index of time.
 $\tau$ Battery storage charging efficiency.
@@
 II Problem Formulation
-Figure 1: DER setup with flexible demands $d\in\mathbb{R}_{+}$, renewable DG $g\in\mathbb{R}_{+}$, storage operation $e\in\mathbb{R}$, and net-consumption $z\in\mathbb{R}$.
-The net energy metering tariff with demand charges is also demonstrated with cost associated with amount of net consuming, net producing and the peak load over the period of day where peak load is recorded.
+Figure 1: DER setup with flexible demand $d\in\mathbb{R}+$, renewable generation $g\in\mathbb{R}+$, storage operation $e\in\mathbb{R}$, and net consumption $z\in\mathbb{R}$.
+The illustrated NEM tariff includes both energy and demand charges, with costs associated with net consumption, net generation, and the peak load recorded over the daily billing period.
 We study a HEMS for prosumers that co-optimizes flexible demand, storage operation, and renewable distributed generation under a NEM tariff with distinct buy/sell rates and peak demand charges.
@@
 As shown in Fig.
-1, the energy storage and flexible loads are jointly scheduled to reduce the demand charge by shaving the load curve, and reduce the volumetric charge by arbitraging NEM rates, while still satisfying prosumer’s consumption.
+1, the energy storage and flexible loads are jointly scheduled to reduce the demand charge by shaving the load curve, and reduce the volumetric charge by arbitraging NEM rates, while still satisfying the prosumer’s consumption.
 II-A Prosumer Resources
@@
 The renewable DG is modeled as an exogenous (positive) Markov random process.
-The flexible demands are denoted by the vector
+The prosumer flexible demands are denoted by the vector
 $\bm{d}_{t}=(d_{t}^{1},\ldots,d_{t}^{K})\in\mathcal{D}\coloneqq\{\bm{d}:\bm{0}\preceq\bm{d}\preceq\overline{\bm{d}}\}\subseteq\mathbb{R}^{K}_{+},$ (1)
-where $K$ is the total number of prosumer devices. Here $\overline{\bm{d}}$ is the consumption bundle’s upper limit.
-[…]
-The prosumer also owns a BTM battery energy storage, which has output $e_{t}\in[-\underline{e},\overline{e}]$, where $\underline{e}$ and $\overline{e}$ are discharging and charging limits.
-The output is decomposed into charging $[e_{t}]^{+}=\max\{e_{t},0\}$ and discharging $[e_{t}]^{-}=\max\{-e_{t},0\}$.
-With charge/discharge efficiencies $\tau,\rho\in(0,1]$, the state of charge (SoC) $s_{t}\in[0,B]$ evolves as:
+where $K$ is the total number of devices, and $\overline{\bm{d}}$ is the consumption bundle’s upper limit.
+[^3] The minimum consumption limit is set to $\bm{0}$ without loss of generality.
+Some important devices might be less flexible or inflexible.
+Mathematically, this is incorporated by assigning these devices stricter flexibility limits, i.e., for an inflexible device $k$, whose consumption should not deviate from $\tilde{d}_{t}^{k}$, its flexibility limits would be $\tilde{d}_{t}^{k}\leq d_{t}^{k}\leq\tilde{d}_{t}^{k}$.
+The prosumer also owns a BESS, which has output $e_{t}\in[-\underline{e},\overline{e}]$, where $\underline{e}$ and $\overline{e}$ are the discharging and charging limits, respectively.
+The output is composed of charging $e_{t}]^{+}=\max\{e_{t},0\}$ and discharging $e_{t}]^{-}=\max\{-e_{t},0\}$ actions.
+Given charging and discharging efficiencies, $\tau\in(0,1]$ and $\rho\in(0,1]$, respectively, the BESS state of charge (SoC) $s_{t}\in[0,B]$ evolves as:
 $s_{t+1}=s_{t}+\tau[e_{t}]^{+}-[e_{t}]^{-}/\rho,\quad t=0,\ldots,T-1,$ (2)
@@
 II-B System Dynamics
-The system state at time $t$, $x_{t}$ is described by $x_{t}=[s_{t},g_{t},c_{t}]$, where $c_{t}$ is the prosumer’s peak demand before time $t$.
-We allow the system to be controlled by battery charging/discharging and consumption adjustments, i.e., the control action is $u_{t}\coloneqq(e_{t},\bm{d}_{t})$.
+The system state at time $t$, $x_{t}$, is described by $x_{t}=[s_{t},g_{t},c_{t}]$, where $c_{t}$ is the prosumer’s peak demand before time $t$.
+The system is controlled by the BESS action, i.e., charging or discharging, and consumption adjustments, hence the control action is $u_{t}\coloneqq(e_{t},\bm{d}_{t})$.
 The peak demand in the state evolves as
@@
 and the battery SoC evolves as per (2).
-II-C NEM X Tariff with Demand Charges
-Under NEM, when the prosumer’s gross loads (from consumption and storage charging) are higher than its generation (from renewable DG), the prosumer pays at the retail rate.
-Otherwise, the prosumer gets compensated at the export rate.
-Let $(p_{t}^{+}\geq p_{t}^{-},p)\geq 0$ be the retail, export, and peak demand prices, respectively.
-We equivalently break down the peak demand charge at the end of time $T$ to each time step $t$ since $pc_{T}=p\sum_{t=0}^{T-1}[z_{t}-c_{t}]^{+}$.
-Thus, we can write the payment (cost) under NEM X in [3] for each time $t$ as
-$P_{t}(z_{t},c_{t})=p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}+p[z_{t}-c_{t}]^{+},$ (5)
-where $P_{t}(\cdot)>0$ ($P_{t}(\cdot)<0$) indicates that the prosumer pays (gets compensated) to (by) the DU.
+II-C NEM Tariff with Demand Charges
+Under NEM, when the prosumer’s gross load (from consumption and storage charging) is higher than its generation (from renewable DG and storage discharging), the prosumer pays at the retail rate.
+Otherwise, the prosumer gets compensated at a predetermined export rate.
+To better reflect the capacity costs incurred by the DU, a demand charge is added to the prosumer payment.
+The demand charge is a one-time payment based on the peak demand over the billing period.
+Let $(p_{t}^{+}\geq p_{t}^{-},p)\geq 0$ be the retail, export, and peak demand prices, respectively. The payment under NEM $P_{t}(\cdot)$ over the considered time horizon is given by
+$\sum_{t=0}^{T-1}P_{t}(z_{t},c_{t})=\sum_{t=0}^{T-1}\left(p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}\right)+pc_{T}+A,$ (5)
+where $A$ denotes a fixed charge, commonly known as a connection charge, that does not vary with consumption [3].
+The payment in (5) corresponds to a three-part tariff structure, consisting of volumetric, demand, and fixed components [28].
+For the convenience of analysis, we equivalently break down the peak demand charge at the end of time $T$ to each time step $t$ as a rolling-based incremental charge, since
+[…]
+Thus, we can write the payment under NEM X in (5) for each time $t$ as
+$P_{t}(z_{t},c_{t})=p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}+p[z_{t}-c_{t}]^{+}+A_{t},$ (6)
+where, for every time $t$, $P_{t}(\cdot)>0$ ($P_{t}(\cdot)<0$) indicates that the prosumer pays (gets compensated) to (by) the DU.
+[^4] We assume $A_{t}=0,\forall t$, since fixed costs do not influence DER scheduling decisions, at least in the short-run.
+Special cases of the general NEM tariff in (6) are considered in [19] when demand charges ($p[z_{t}-c_{t}]^{+}$) are absent and in [4] when energy exports are absent ($p_{t}^{-}[z_{t}]^{-}$).
 II-D Prosumer Utility Function
@@
 II-E DER Co-Optimization Problem
-To optimally schedule its DER, the prosumer solves a dynamic program that maximizes its reward function by explicitly accounting for the temporal coupling of storage operations and an NEM tariff with bidirectional energy pricing and peak demand charges.
+To optimally schedule its DER, the prosumer solves a dynamic program that maximizes its reward function by explicitly accounting for the temporal coupling of storage operations and a NEM tariff with asymmetric retail and export rates and peak demand charges.
 The reward function is defined as the difference between the utility of consumption (7) and payment under NEM and demand charges (6). Therefore, the reward function is
-$r_{t}(x_{t},u_{t})=r_{t}((s_{t},g_{t},c_{t}),(e_{t},\mathbf{d}_{t}))=U_{t}(\bm{d}_{t})-P_{t}(z_{t},c_{t}),\quad t=0,\ldots,T-1.$ (7)
+$r_{t}(x_{t},u_{t})$ $=r_{t}((s_{t},g_{t},c_{t}),(e_{t},\mathbf{d}_{t}))$ (8)
+$=U_{t}(\bm{d}_{t})-P_{t}(z_{t},c_{t}),\quad t=0,\ldots,T-1.$
 To encourage the preservation of energy in the battery, we define a linear reward to the remaining energy at the terminal stage $T$, i.e.
@@
 where $\gamma>0$ is the value assigned to the remaining charge in the storage at the end of the horizon.
-The policy $\pi=(\pi_{0},\ldots,\pi_{T-1})$ is a sequence of mapping from state space to action space such that $\pi_{t}(x_{t})=u_{t}$.
+The policy $\bm{\pi}=(\pi_{0},\ldots,\pi_{T-1})$ is a sequence of mappings from state space to action space such that $\pi_{t}(x_{t})=u_{t}$, for every $t$.
 Given initial state $x_{0}=(s,g,0)$, the storage-consumption co-optimization is then defined as:
 $\max_{\pi}\quad$ $\mathbb{E}\left[\sum_{t=0}^{T}r_{t}(x_{t},u_{t})\right]$ (10a)
-Subject to for $\quad\text{for all}\ t=0,1,\ldots,T-1$
+Subject to $\quad\text{for all}\ t=0,1,\ldots,T-1$
 $x_{0}=(s,g,0)$ (10b)
@@
 $\bm{0}\preceq\bm{d}_{t}\preceq\bm{\overline{d}}.$ (10i)
-Constraints (9e) and (9h) induce intertemporal coupling, precluding a myopic (stage-wise optimal) policy.
-Specifically, relaxing either constraint in isolation is insufficient to eliminate the temporal dependencies, and the resulting optimal policy remains non-myopic (see Thm.
-4 in the Appendix).
-However, when both constraints are simultaneously relaxed, the problem admits a myopic optimal policy characterized by decoupled per-period decisions (see Lem.
-4).
+Constraints (10e) and (10h) induce intertemporal coupling, precluding a myopic (stage-wise) optimal policy.
+Specifically, relaxing either constraint in isolation is insufficient to eliminate the temporal dependencies, and the resulting optimal policy remains non-myopic (see Theorem 4 in Appendix -A).
+However, when both constraints are simultaneously relaxed, the problem admits a myopic optimal policy characterized by decoupled per-period decisions (see Lemma 4).
 The value function $V_{t}^{\pi}(\cdot)$ is defined as the expected total reward received under the policy $\pi$, i.e., $V_{t}^{\pi}(x)=\mathbb{E}[\sum_{t=0}^{T}r_{t}(x_{t},\pi(x_{t}))|x_{t}=x]$.
@@
 III Dynamic Program Analytical Results
-[…]
-[…]
-The following lemma establishes a conclusion regarding the concavity of the optimal value function.
+In this section, we show that the optimal policy for the co-optimization problem in §II-E is a threshold-based one.
+First, we show that the optimal value function is concave.
 Lemma 1.
@@
 We follow the proof technique for Lemma 1 in [4].
-[…]
-[…]
-[…]
-First, by the concavity of $U_{t}$:By definition, $U_{t}$ is concave:
+Consider two different states $x_{\tau}$ and $x_{\tau}^{\prime}$ with their corresponding optimal trajectories $\{(x_{t},u_{t})\}_{t=\tau}^{T}$ and $\{(x_{t}^{\prime},u_{t}^{\prime})\}_{t=\tau}^{T}$.
+Let $\tilde{x}_{\tau}=(x_{\tau}+x_{\tau}^{\prime})/2$ be their average initial state, and define the average policy $\tilde{\pi}=\{\tilde{u}_{t}=(u_{t}+u_{t}^{\prime})/2\}_{t=\tau}^{T}$ with corresponding trajectory $\{\tilde{x}_{t}\}_{t=\tau}^{T}$.
+First, by the concavity of $U_{t}$:
 $(U_{t}(\bm{d}_{t})+U_{t}(\bm{d}_{t}^{\prime}))/2\leq U_{t}((\bm{d}_{t}+\bm{d}_{t}^{\prime})/{2})=U_{t}(\tilde{\bm{d}}_{t}).$ (11)
-[…]
 The cost $P_{t}(z_{t},c_{t})$ consists of two components: the energy cost $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ and the demand charging $p[z_{t}-c_{t}]^{+}$.
-Since $p_{t}^{+}\geq p_{t}^{-}\geq 0$, the energy cost is convex.We separate the $P_{t}(z_{t},c_{t})$ into energy cost $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ and the demand charging cost $p[z_{t}-c_{t}]^{+}$.
-By assumption, $p_{t}^{+}\geq p_{t}^{-}\geq 0$, and $p_{t}^{+}[0]^{+}=p_{t}^{-}[0]^{-}=0$, the left derivative at 00 is smaller than the right derivative at 00.
-So, the derivative of $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ is non-decreasing and hence $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ is convex.
+Since $p_{t}^{+}\geq p_{t}^{-}\geq 0$, the energy cost is convex.
 Then,
-$-\sum_{t=\tau}^{T}(p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}+p_{t}^{+}[z_{t}^{\prime}]^{+}-p_{t}^{-}[z_{t}^{\prime}]^{-})/2$ $\leq-\sum_{t=\tau}^{T}p_{t}^{+}[(z_{t}+z_{t}^{\prime})/2]^{+}-p_{t}^{-}[(z_{t}+z_{t}^{\prime})/2]^{-}$ (11)
+$-\sum_{t=\tau}^{T}(p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}+p_{t}^{+}[z_{t}^{\prime}]^{+}-p_{t}^{-}[z_{t}^{\prime}]^{-})/2$ (12)
+$\leq-\sum_{t=\tau}^{T}p_{t}^{+}[(z_{t}+z_{t}^{\prime})/2]^{+}-p_{t}^{-}[(z_{t}+z_{t}^{\prime})/2]^{-}$
 $=-\sum_{t=\tau}^{T}p_{t}^{+}[\tilde{z}_{t}]^{+}-p_{t}^{-}[\tilde{z}_{t}]^{-},$
 and the demand charging cost becomes
-$-\sum_{t=\tau}^{T}(p[z_{t}-c_{t}]^{+}+p[z_{t}^{\prime}-c_{t}^{\prime}]^{+})/2=-p(\max_{\tau\leq t\leq T}\{z_{t}\}-c_{\tau}+\max_{\tau\leq t\leq T}\{z_{t}^{\prime}\}-c_{\tau}^{\prime})/2$ $\leq-p(\max_{\tau\leq t\leq T}\{(z_{t}+z_{t}^{\prime})/2\}-(c_{\tau}+c_{\tau}^{\prime})/2)$ (12)
+$-\sum_{t=\tau}^{T}(p[z_{t}-c_{t}]^{+}+p[z_{t}^{\prime}-c_{t}^{\prime}]^{+})/2$ (13)
+$=$ $-p(\max_{\tau\leq t\leq T}\{z_{t}\}-c_{\tau}+\max_{\tau\leq t\leq T}\{z_{t}^{\prime}\}-c_{\tau}^{\prime})/2$
+$\leq$ $-p(\max_{\tau\leq t\leq T}\{(z_{t}+z_{t}^{\prime})/2\}-(c_{\tau}+c_{\tau}^{\prime})/2)$
 $=$ $-p(\max_{\tau\leq t\leq T}\{\tilde{z}_{t}\}-\tilde{c}_{\tau})=\sum_{t=\tau}^{T}p[\tilde{z}_{t}-\tilde{c}_{t}]^{+}.$
@@
 which proves the concavity of the optimal value function. ∎
-[…]
 It turns out that the optimal value function is also non-decreasing in the system state, which is formally stated in the next lemma.
 Lemma 2.
-[…]
+The optimal value function $V_{t}^{*}(s_{t},g_{t},c_{t})$ is non-decreasing in $s_{t}$, $g_{t}$, and $c_{t}$ for every time step $t\leq T$.
 Proof.
@@
 (3) $c_{t}$: From (6), $P_{t}(z_{t},c_{t})$ is non-increasing in $c_{t}$, hence $V_{t}^{*}(s_{t},g_{t},c_{t})$ is non-decreasing in $c_{t}$. ∎
-[…]
 Lastly, we derive a property that links the optimal $Q$-function, defined as $Q_{t}^{*}(x_{t},u_{t})=r(x_{t},u_{t})+\mathbb{E}[V_{t+1}^{*}(x_{t+1})]$, to the system’s control action $u_{t}$.
@@
 Therefore, $Q_{t}^{*}(x_{t},u_{t})$ is concave in $u_{t}$. ∎
-Building on the results of Lem. 1, Lem. 2, and Lem. 3, we establish Thm. 1, which characterizes the structure of the optimal co-optimization policy under demand charges.
+Building on the results of Lemma 1, Lemma 2, and Lemma 3, we establish Theorem 1, which describes the structure of the optimal co-optimization policy under demand charges.
 Theorem 1.
@@
 Proof.
-[…]
+The feasible region for action $u_{t}$, $(e_{t},\bm{d}_{t})\in\mathcal{F}_{t}=[\max\{-\underline{e},-\rho s_{t}\},\min\{\overline{e},(B-s_{t})/\tau\}]\times\prod_{i=1}^{K}[0,d_{i}]$ is a compact set.
+From Lemma 1 to Lemma 3, the optimal value function and $Q$-function is concave in action $u_{t}$, so the optimal action
 $u_{t}^{*}=\arg\max_{u\in\mathcal{F}_{t}}\{r_{t}(x_{t},u)+Q_{t}^{*}(x_{t},u)\}$ (15)
-[…]
 must exist in the interior of the feasible set $\mathcal{F}_{t}$ or at its boundary.
 The optimal action can be obtained by solving concave optimizations with and without boundary constraints, which results in a threshold policy.
-So, we can get the optimal solution by solving $(K+1)2^{K}$ constrained concave optimization and one unconstrained concave optimization, which results in a threshold solution of $(K+1)2^{K}+1$ cases.
 ∎
-Although Thm.
-1 reveals a structured policy, direct dynamic programming implementation remains impractical.
-First, optimization necessitates gradients of $r_{t}$ and $Q_{t}$, which are computationally prohibitive for complex utility functions.
-Also, continuous-space Q-function tracking requires costly discretization, as validated in simpler settings [4].
-To this end, we develop an efficient approximation algorithm to overcome these computational barriers.
-IV Efficient Approximation Algorithm
-Since computing the exact optimal solution for problem (9) is computationally intractable, we propose an efficient approximation algorithm.
-The large storage peak searching (LSPS) algorithm efficiently searches for the optimal peak demand by solving a relaxed version of the original problem.
+Although Theorem 1 reveals a structured policy, direct dynamic programming implementation remains impractical.
+If we directly apply concave optimization on the compact sets, it necessitates gradients of $r_{t}$ and $Q_{t}$, which are computationally prohibitive, especially for complex utility functions that may not be additive with respect to consuming devices.
+On the other hand, if we want to avoid the expensive optimization by backtracking, the continuous-space $Q$-function tracking requires costly discretization, as validated in simpler settings in [4].
+To this end, we develop, in §IV, an efficient approximation algorithm to overcome these computational barriers.
+IV An Efficient Approximation Algorithm
+Since obtaining the exact optimal solution to problem (10) is computationally intractable, we develop an efficient approximation algorithm.
+The proposed large storage peak searching (LSPS) algorithm efficiently searches for the optimal solution by solving a relaxed version of the original problem.
 The resulting solution is then projected onto the feasible region $\mathcal{F}_{t}$, producing near-optimal and computationally tractable approximations.
 The original formulation prevents myopic solutions through constraints (10e) and (10h).
-To address this computational challenge, we first relax constraint (9e) by considering a large capacity regime (Sec.
-IV-A), then develop an efficient method to handle constraint (9h) (Sec.
-IV-B).
+To address this computational challenge, we first relax constraint (10e) by considering a large storage capacity regime (see §IV-A), then develop an efficient method to handle constraint (10h) (see §IV-B).
 IV-A Relaxation of the DER Co-Optimization Problem
-We consider a large capacity regime as a relaxation of the battery capacity constraint (9e).
-Observing that $e_{t}$ and $\mathbf{d}_{t}$ consistently appear together in the expression $\mathbf{1}^{\top}\mathbf{d}_{t}+e_{t}$, we introduce a helper function $h_{t}(v_{t})$:
-$h_{t}(v_{t}):=\max\{U_{t}(\bm{d}_{t})+\gamma e_{t}:\bm{1}^{\top}\bm{d}_{t}+e_{t}=v_{t},\bm{0}\preceq\bm{d}_{t}\preceq\overline{\bm{d}},-\underline{e}\leq e_{t}\leq\overline{e}\},$ (15)
+We consider a large storage capacity regime as a relaxation of the battery capacity constraint (10e).
+Observing that $e_{t}$ and $\bm{d}_{t}$ consistently appear together in the expression $\mathbf{1}^{\top}\bm{d}_{t}+e_{t}$, we introduce a helper function $h_{t}(v_{t})$:
+$h_{t}(v_{t})=\max\{$ $U_{t}(\bm{d}_{t})+\gamma e_{t}:\bm{1}^{\top}\bm{d}_{t}+e_{t}=v_{t},$ (16)
+$\bm{0}\preceq\bm{d}_{t}\preceq\overline{\bm{d}},-\underline{e}\leq e_{t}\leq\overline{e}\},$
 where $v_{t}$ represents the net consumption excluding renewables at time step $t$.
 The function $h_{t}(v_{t})$ is differentiable, monotonically increasing, and concave for all time steps $t$.
-For any $v_{t}\in[-\underline{e},\overline{e}+\bm{1}^{\top}\bm{\overline{d}}]$, there exists a valid pair $(e_{t},\mathbf{d}_{t})$ that achieves $h_{t}(v_{t})$ with $\mathbf{1}^{\top}\mathbf{d}_{t}+e_{t}=v_{t}$.
+For any $v_{t}\in[-\underline{e},\overline{e}+\bm{1}^{\top}\bm{\overline{d}}]$, there exists a valid pair $(e_{t},\bm{d}_{t})$ that achieves $h_{t}(v_{t})$ with $\mathbf{1}^{\top}\bm{d}_{t}+e_{t}=v_{t}$.
 This equivalence allows us to optimize over $v_{t}$ after relaxing constraint (10e).
-Using the helper function (15), we formulate the relaxed problem as:
-$\max_{v_{0},\ldots,v_{T-1}}\quad$ $\sum_{t=0}^{T-1}h_{t}(v_{t})-p_{t}^{+}[v_{t}-g_{t}]^{+}+p_{t}^{-}[v_{t}-g_{t}]^{-}-p[v_{t}-g_{t}-c_{t}]^{+}$
-Subject to $-\underline{e}\leq v_{t}\leq\overline{e}+\bm{1}^{\top}\overline{\bm{d}}$ (16a)
-$c_{t+1}=\max\{c_{t},v_{t}-g_{t}\},$ (16b)
-To eliminate the temporal dependencies introduced by constraint (9h), we consider a fixed peak demand bound $c$ in the large capacity regime (16):
+Using the helper function (16), we formulate the relaxed problem as
+$\max_{v_{0},\ldots,v_{T-1}}\quad$ $\sum_{t=0}^{T-1}h_{t}(v_{t})-p_{t}^{+}[v_{t}-g_{t}]^{+}+p_{t}^{-}[v_{t}-g_{t}]^{-}$
+$-p[v_{t}-g_{t}-c_{t}]^{+}$ (17a)
+Subject to $-\underline{e}\leq v_{t}\leq\overline{e}+\bm{1}^{\top}\overline{\bm{d}}$ (17b)
+$c_{t+1}=\max\{c_{t},v_{t}-g_{t}\}.$ (17c)
+To eliminate the temporal dependencies introduced by constraint (10h), we consider a fixed peak demand bound $c$ in the large capacity regime (17)
 $\max_{v_{0},\ldots,v_{T-1}}\quad$ $\sum_{t=0}^{T-1}h_{t}(v_{t})-p_{t}^{+}[v_{t}-g_{t}]^{+}+p_{t}^{-}[v_{t}-g_{t}]^{-}-pc$ (18a)
@@
 $v_{t}\leq g_{t}+c.$ (18c)
-This reformulation replaces the recursive peak demand update (16b) with a single-step upper-bound constraint (17c), effectively removing temporal dependencies while preserving the optimal policy structure.
+This reformulation replaces the recursive peak demand update (17c) with a single-step upper-bound constraint (18c), effectively removing temporal dependencies while preserving the optimal policy structure.
 We show that the optimal policy of (17) is preserved under this relaxation.
@@
 ∎
-By Thm. 2, we can compute the optimal policy for (16) by searching for the optimal $c^{*}$ and solving (17).
-IV-B Solution under Given Peak Demand
+By Theorem 2, we can compute the optimal policy for (17) by searching for the optimal $c^{*}$ and solving (18).
+IV-B Solution for a Given Peak Demand
 With a fixed peak demand bound $c$, the optimization problem decomposes into independent single-step subproblems that can be solved analytically.
@@
 For $t=0,\ldots,T-1$, the optimal policy for (18) is given by
-$v_{t}^{\ast}(g_{t})=\begin{cases}-\underline{e},\quad\quad\quad\quad\quad\ v^{\dagger}(g_{t})\leq-\underline{e},\\ v_{t}^{\dagger}(g_{t}),\quad\quad\quad\quad-\underline{e}\leq v^{\dagger}(g_{t})\leq\min\{c+g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}\},\\ \min\{c+g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}\},\min\{c+g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}\}\leq v^{\dagger}(g_{t}),\end{cases}$ (18)
-where the unconstrained optimum $v_{t}^{\dagger}(g_{t})$ is the solution without constraints (17b) and (17c), given by
-$v_{t}^{\dagger}(g_{t})=\begin{cases}(h^{\prime})^{-1}(p_{t}^{-}),&(h^{\prime})^{-1}(p_{t}^{-})<g_{t},\\ g_{t},&(h^{\prime})^{-1}(p_{t}^{+})\leq g_{t}\leq(h^{\prime})^{-1}(p_{t}^{-}),\\ (h^{\prime})^{-1}(p_{t}^{+}),&g_{t}<(h^{\prime})^{-1}(p_{t}^{+}).\end{cases}$ (19)
+$v_{t}^{\ast}(g_{t})=$
+$\begin{cases}-\underline{e},&\text{if}\quad v^{\dagger}_{t}(g_{t})<-\underline{e},\\ v_{t}^{\dagger}(g_{t}),&\text{if}\quad-\underline{e}\leq v^{\dagger}_{t}(g_{t})\leq\Gamma_{t},\\ \Gamma_{t},&\text{if}\quad\Gamma_{t}<v^{\dagger}_{t}(g_{t}),\end{cases}$ (19)
+where $\Gamma_{t}=\min\{c+g_{t},\overline{e}+\bm{1}^{\top}\overline{\bm{d}}\}$ is the upper bound of $v_{t}^{*}(g_{t})$, the unconstrained optimum $v_{t}^{\dagger}(g_{t})$ is the solution without constraints (18b) and (18c), given by
+$v_{t}^{\dagger}(g_{t})$
+$=$ $\begin{cases}(h^{\prime})^{-1}(p_{t}^{-}),&\text{if}\quad(h^{\prime})^{-1}(p_{t}^{-})<g_{t},\\ g_{t},&\text{if}\quad(h^{\prime})^{-1}(p_{t}^{+})\leq g_{t}\leq(h^{\prime})^{-1}(p_{t}^{-}),\\ (h^{\prime})^{-1}(p_{t}^{+}),&\text{if}\quad g_{t}<(h^{\prime})^{-1}(p_{t}^{+}).\end{cases}$ (20)
 Proof.
-With fixed peak demand, the optimization decomposes into independent subproblems at each time step $t=0,1,\cdots,T-1$:
+With fixed peak demand, the optimization decomposes into independent subproblems at each time step $t=0,1,\ldots,T-1$:
 $\max_{v_{t}}\quad$ $h_{t}(v_{t})-p_{t}^{+}[v_{t}-g_{t}]^{+}+p_{t}^{-}[v_{t}-g_{t}]^{-}$ (21a)
@@
 The unconstrained optimum is
-$v_{t}^{\dagger}(g_{t})=\arg\max_{v}~{}~{}h_{t}(v)-p_{t}^{+}[v-g_{t}]^{+}+p_{t}^{-}[v-g_{t}]^{-}=\begin{cases}(h_{t}^{\prime})^{-1}(p_{t}^{-}),&(h_{t}^{\prime})^{-1}(p_{t}^{-})<g_{t},\\ g_{t},&(h_{t}^{\prime})^{-1}(p_{t}^{+})<g_{t}<(h_{t}^{\prime})^{-1}(p_{t}^{-}),\\ (h_{t}^{\prime})^{-1}(p_{t}^{+}),&g_{t}<(h_{t}^{\prime})^{-1}(p_{t}^{+}).\end{cases}$
+$v_{t}^{\dagger}(g_{t})=\arg\max_{v}h_{t}(v)-p_{t}^{+}[v-g_{t}]^{+}+p_{t}^{-}[v-g_{t}]^{-}$
+[…]
 Projecting onto the feasible region defined by constraint (21b) yields the optimal policy in (4). ∎
-The optimal policy has a clear economic interpretation based on the relationship between renewable generation and consumption needs.
-The unconstrained solution $v_{t}^{\dagger}(g_{t})$ balances the marginal utility of consumption against energy trading costs.
-When $(h_{t}^{\prime})^{-1}(p_{t}^{-})<g_{t}$, the renewable generation $g_{t}$ exceeds the energy demanded by consuming at the selling price $\pi^{-}$ and charging the battery.
-In this case, the optimal consumption $v_{t}^{\dagger}(g_{t})=(h_{t}^{\prime})^{-1}(p_{t}^{-})$ is set such that the marginal utility equals the selling price, meaning excess generation beyond this consumption level should be sold to the grid.
-When $(h_{t}^{\prime})^{-1}(p_{t}^{+})\leq g_{t}\leq(h_{t}^{\prime})^{-1}(p_{t}^{-})$, the renewable generation exactly matches the optimal consumption level, so $v_{t}^{\dagger}(g_{t})=g_{t}$ and no energy trading occurs.
+The optimal policy has an intuitive economic interpretation based on the relationship between renewable generation and consumption needs.
+The unconstrained optimal solution $v_{t}^{\dagger}(g_{t})$ balances the marginal utility of consumption against energy costs.
+When $(h_{t}^{\prime})^{-1}(p_{t}^{-})<g_{t}$, the renewable generation $g_{t}$ exceeds the energy demanded by consuming at the export price $p^{-}$ and charging the battery.
+In this case, the optimal consumption $v_{t}^{\dagger}(g_{t})=(h_{t}^{\prime})^{-1}(p_{t}^{-})$ is set such that the marginal utility equals the export price, meaning excess generation beyond this consumption level should be sold to the grid.
+When $(h_{t}^{\prime})^{-1}(p_{t}^{+})\leq g_{t}\leq(h_{t}^{\prime})^{-1}(p_{t}^{-})$, the renewable generation exactly matches the optimal consumption level, so $v_{t}^{\dagger}(g_{t})=g_{t}$ and no energy transactions occur.
 When $g_{t}<(h_{t}^{\prime})^{-1}(p_{t}^{+})$, renewable generation is insufficient relative to the consumption level that would justify purchasing energy at price $p_{t}^{+}$.
@@
 The final policy $v_{t}^{*}(g_{t})$ then projects this unconstrained optimum onto the feasible region, ensuring that consumption respects both battery operation limits and peak demand constraints.
-IV-C Algorithmic Search of the Optimal Peak Demand
+IV-C Algorithmic Search for the Optimal Peak Demand
 We leverage the concavity properties of the objective function to efficiently determine the optimal peak demand $c^{*}$ without solving the complete optimization problem (17).
@@
 Let $J(c)$ be the optimal objective function of (18),
-$J(c):=\sum_{t=0}^{T-1}h_{t}(v_{t}^{*})-p_{t}^{+}[v_{t}^{*}-g_{t}]^{+}+p_{t}^{-}[v_{t}^{*}-g_{t}]^{-}-pc,$ (21)
+$J(c):=-pc+\sum_{t=0}^{T-1}h_{t}(v_{t}^{*})-p_{t}^{+}[v_{t}^{*}-g_{t}]^{+}+p_{t}^{-}[v_{t}^{*}-g_{t}]^{-},$ (22)
 where $v_{t}^{*}$ depends on $c$ through the optimization. Then $J(c)$ is concave in $c$.
@@
 Let $I(c)=\{t:v_{t}^{*}=c+g_{t}\}$ be the set of time indices where the peak demand constraint is active. We can decompose the objective function as
-$J(c)=-pc+\sum_{t\in I(c)}h_{t}(c+g_{t})-p_{t}^{+}c+\sum_{t\in[T-1]\setminus I(c)}h_{t}(v_{t}^{*})-p_{t}^{+}[v_{t}^{*}-g_{t}]^{+}+p_{t}^{-}[v_{t}^{*}-g_{t}]^{-}.$
+$J(c)=$ $-pc+\sum_{t\in I(c)}h_{t}(c+g_{t})-p_{t}^{+}c$
+$+\sum_{t\in[T-1]\setminus I(c)}h_{t}(v_{t}^{*})-p_{t}^{+}[v_{t}^{*}-g_{t}]^{+}+p_{t}^{-}[v_{t}^{*}-g_{t}]^{-}.$
 As $c$ increases, the constraint set becomes less restrictive, so $|I(c)|$ is non-increasing.
@@
 IV-D Large Storage Peak Searching (LSPS) Algorithm
-The LSPS algorithm, in Alg.
-1, efficiently approximates the problem (9) by combining the relaxation techniques from the previous sections.
-It leverages the key structural properties established in the previous sections: the decomposition enabled by the large capacity relaxation, the analytical solution for fixed peak demand, and the concavity of the objective function in the peak demand parameter.
+The LSPS algorithm (Algorithm 1) is designed to solve the scheduling problem offline.
+For a given scheduling horizon (e.g., 24 hours), the algorithm uses forecasts of exogenous variables, like renewable generation $g_{t}$, for the entire period to solve for a single optimal peak demand, $c^{*}$.
+This single $c^{*}$ is then used to determine the optimal control actions for each time step within the horizon.
+The algorithm efficiently approximates the problem (10) by combining the relaxation techniques from the previous sections.
+It leverages the key structural properties established in the previous sections: the decomposition enabled by the large storage capacity relaxation, the analytical solution for fixed peak demand, and the concavity of the objective function in the peak demand parameter.
 By searching over the discrete candidate set $\mathcal{C}$, we avoid solving the full dynamic optimization problem and rather solve a linear complexity algorithm, as formalized below, while maintaining a small solution gap compared to the optimal.
@@
 2:Derivatives $J^{\prime}(\mathcal{C})\leftarrow\{J^{\prime}(c):c\in\mathcal{C}\}$
-3:$c_{1}\leftarrow\max\{c:c\in\mathcal{C},J^{\prime}(c)<0\}$
+[…]
 4:Solve $J^{\prime}(c^{*})=0$ using (24) for optimal peak demand $c^{*}$
 5:Find optimal actions $v_{t}^{*}$ using (4) with peak demand $c^{*}$
-6:Project $v_{t}^{*}$ back to variables $(e_{t}^{*},\mathbf{d}_{t}^{*})$ using (15)
+6:Project $v_{t}^{*}$ back to variables $(e_{t}^{*},\bm{d}_{t}^{*})$ using (16)
 7:Clip $e_{t}^{*}$ to ensure $s_{t+1}\in[0,B]$
 Theorem 3.
-Alg. 1 has computational complexity $\mathcal{O}(T)$.
+Algorithm 1 has computational complexity $\mathcal{O}(T)$.
 Proof.
@@
 ∎
-Empirical analysis on the performance of the LSPS algorithm is presented in Sec. V.
+Empirical analysis on the performance of the LSPS algorithm is presented in §V.
+A comparison between the computational cost of the LSPS algorithm and a deterministic optimization that solves (10) offline in one-shot is presented in Appendix -C.
 V Case Study
-Using two rich datasets, and considering a one-day operation horizon $T=24$, we evaluate the DER co-optimization performance of the proposed LSPS algorithm and an RL controller using a Proximal Policy Optimization (PPO) algorithm compared to a theoretical maximum that assumes perfect knowledge of future renewables solving Eq.
-9.
-The performance is measured by the solution gap, which is the difference between the cumulate reward of the algorithms and the theoretical upper bound.
-Experiments were run on a server with 1 AMD Ryzen Threadripper 3990X 64-Core Processor and 4 Nvidia RTX A4000 GPUs.
+Using two rich datasets, and considering a one-day operation horizon with hourly decisions, $T=24$ hours, we evaluate the DER co-optimization performance of the proposed LSPS algorithm and several benchmarks compared to a theoretical maximum that assumes perfect knowledge of future renewables solving (10).
+Performance is measured by the solution gap, which is the difference between the cumulative reward of the algorithms and the theoretical upper bound.
+The experiments were run on a server with 1 AMD Ryzen Threadripper 3990X 64-Core Processor and 4 Nvidia RTX A4000 GPUs.
 V-A Dataset Description
-(a) Daily net demand for HouseZero (Jun 2022 – May 2024)
-(b) Daily net demand for BDGP building (Jan 2016 – Dec 2017)
+(a) Daily net demand data for HouseZero from Jun 2022 to May 2024
+(b) Daily net demand data for BDGP building from Jan 2016 to Dec 2017
 Figure 2: Daily net demand distributions for HouseZero and BDGP over the training and testing period.
 Two open-source datasets from Harvard HouseZero [30] and Building Data Genome Project (BDGP) [31] were used to evaluate the performance of DER scheduling under different algorithms.
-The two buildings were chosen with distinct characteristics while both are equipped with solar PV.
-Renewable generation and demand data, resampled to hourly granularity, are used for training and testing the performance of different control strategies.
+The two buildings were chosen with distinct characteristics, while both are equipped with solar PV.
+Renewable generation and demand data, resampled to hourly granularity, were used to train and test the performance of different control strategies.
 Fig.
-2 depicts the net demand distribution for HouseZero and BDGP buildings.
+2 shows the net demand distribution for the HouseZero and BDGP buildings.
 The two buildings differ significantly in their energy pattern: HouseZero is a low-energy, ultra-efficient building with a peak demand below 2 kW and a frequently negative net consumption.
@@
 We adopt the concave utility function from [32]: $U_{t}(d_{t})=\alpha_{t}d_{t}-\frac{1}{2}\beta_{t}d_{t}^{2}$, where $\alpha_{t}$ and $\beta_{t}$are time-varying and learned a priori using historical energy prices, demands, and demand elasticity [26].
-The elasticity was set to -0.1 [28].
-To get more crisp insights, we assumed time-invariant NEM parameters, with a $p^{+}=\$0.12kWh$ retail (import) rate and a $p^{-}=\$0.06/kWh$ sell (export) rate.
-The daily peak demand charge was set to $p=\$10/kW$.
-The final salvage value for the remaining energy in the battery was set at $\gamma=\$0.09/kWh$.
-The baseline battery parameters were selected based on the demand data and size of the building.
+The price elasticity of consumption was set to 0.1-0.1 [33].
+To get more crisp insights, we assumed time-invariant NEM tariff parameters, with a retail (import) rate of $p^{+}=\mathdollar 0.12$/kWh and a sell (export) rate of $p^{-}=\mathdollar 0.06$/kWh.
+The daily peak demand charge was set to $p=\mathdollar 10$/kW.
+The final salvage value for the remaining energy in the battery was set at $\gamma=\mathdollar 0.09$/kWh.
+The baseline battery parameters were selected based on the demand data and the size of the building.
 For HouseZero, $B=5$ kWh and $\underline{e}=\overline{e}=1$ kW were used; while for BDGP building, $B=350$ kWh and $\underline{e}=\overline{e}=50$ kW were used.
@@
 V-C Control Strategies
-We present five control strategies to schedule DER here:
+We consider and compare five distinct control strategies for scheduling DER.
+The description of each strategy is provided below, while a detailed comparison of their respective advantages and limitations is presented in Appendix -D.
 Backup Mode
@@
 The battery discharges only during grid outages, equivalent to the case without a battery system [4] during normal operations.
+Consequently, this control strategy cannot exploit the full potential of jointly optimizing demand and storage.
 Renewable-Adjusted Threshold Policy (RATP)
@@
 However, RATP’s performance significantly degrades under NEM with demand charges, as it is agnostic to strategic net consumption decisions that shape demand peaks.
-Mathematically, the storage control under RATP is
-$e_{t}(g_{t})=\begin{cases}\max\{-\underline{e},\rho s_{t},-\tilde{d}_{t}\},&\bar{d}_{t}-g_{t}>0\\ \min\{\overline{e},(B-s_{t})/\tau,-\tilde{d}_{t}\}&\bar{d}_{t}-g_{t}\leq 0,\end{cases}$
-and $d_{t}=\bar{d}_{t}$ in both cases. We use RATP and Threshold interchangeably.
+Mathematically, the storage control under RATP, for a given demand $d_{t}=\hat{d}_{t}$, is
+$e_{t}(g_{t})=\begin{cases}\max\{-\underline{e},-\rho s_{t},-\tilde{d}_{t}\},&\tilde{d}_{t}(g)>0\\ \min\{\overline{e},(B-s_{t})/\tau,-\tilde{d}_{t}\}&\tilde{d}_{t}(g)\leq 0,\end{cases}$
+for $t=0,1,\ldots,T-1$.
 RL Algorithm
@@
 We train the networks with a learning rate of $1e^{-4}$, batch size of 64, and 16 epochs per update.
-Other training parameters are kept as default provided by Stable-baselines3 [29].
+Other training parameters are kept as the default provided by Stable-baselines3 [34].
+Training histories for RL are provided in Appendix -E, showing the convergence of the trained policy that is used in §V-D.
 Large Storage Peak Searching (LSPS) Algorithm
-LSPS algorithm, described in Alg.
-1, offers an efficient and scalable method to co-optimize DER.
-Following Alg.
-1, we first calculate the optimal peak, then, by solving Eq.
-17 with the optimal peak, we get the battery and demand actions at each time step.
+LSPS algorithm, described in Algorithm 1, offers an efficient and scalable method to co-optimize DER.
+Following Algorithm 1, we first calculate the optimal peak, then, by using Lemma 4 with the optimal peak, we compute the co-optimized battery and demand actions at each time step.
 Theoretical Upper Bound
-The theoretical upper bound assumes perfect knowledge of the renewables for the entire scheduling period and solves the constrained convex program in Eq. 9.
+The theoretical upper bound assumes perfect knowledge of the renewables for the entire scheduling period and deterministically solves the constrained convex program in (10) in one-shot while relaxing the simultaneous charging-discharging constraint, giving it a slight advantage over LSPS.
+During the prediction stage, RL takes in the current state values, including battery SoC, renewable generation, peak demand, and outputs the optimal actions based on the learned policy.
+While the LSPS algorithm during prediction takes in the optimal peak demand in addition to the current demand and generation.
+The optimal actions are generated by solving an optimization problem.
+During the training stage, a similar level of computation is required as RL learns the system dynamics, including the patterns of the renewable profile with two years of hour-level training data, while LSPS requires models for renewable predictions.
 V-D Simulation Results
-We compare the performance of the control strategies (a)-(d) above to the theoretical optimal algorithm in (e) by varying the following parameters:
+We compare the performance of the control strategies (a)-(d) above to the theoretical upper bound in (e) by varying the following parameters:
   • Battery capacity ($B$): 5 kWh – 50 kWh for HouseZero and 350 kWh – 950 kWh for BDGP building to simulate battery sizes from small to large.
 […]
-  • Export rate of electricity ($p^{-}$): \$$\$0/\text{kWh}$\$$\$0.12/\text{kWh}$ from no to high sell back incentive with $p^{-}=p^{+}$.
-  • Peak demand price of electricity ($p$): \$$\$0/\text{kW}$\$$\$10/\text{kW}$ from no penalty on a high peak to a reasonably high peak charge commonly practiced in electricity tariffs.
-Tab.
-III summarizes the overall performance comparisons in terms of the solution gap to the optimal solutions.
-[^3] Table IV in the Appendix details all cases for reference.
-Overall, the LSPS algorithm performed closest to the optimal solutions over all scenarios with an average gap of 13.77%, followed by RL of 20.21%, threshold algorithm of 41.93%, and backup mode of 43.75%.
+  • Export rate of electricity ($p^{-}$): $\mathdollar 0/\text{kWh}$$\mathdollar 0.12/\text{kWh}$ from no to high sell back incentive with $p^{-}=p^{+}$.
+  • Peak demand price of electricity ($p$): $\mathdollar 0/\text{kW}$$\mathdollar 10/\text{kW}$ from no penalty on a high peak to a reasonably high peak charge commonly practiced in electricity tariffs.
+Table III summarizes the overall performance comparisons in terms of the solution gap to the optimal solutions.
+[^5] Table V in Appendix -B provides detailed results for all cases.
+Overall, the LSPS algorithm performed closest to the optimal solutions over all scenarios with an average gap of 12.86%, followed by RL of 20.21%, RATP algorithm of 41.93%, and backup mode of 43.75%.
 The LSPS algorithm was closest to the optimal 34 out of 42 scenarios (81%) while RL was closest to optimal in the remaining 8 scenarios (19%).
-Furthermore, for almost all cases, control strategies/algorithms showed better performance for the smaller building of HouseZero, with LSPS performing only 4.52% below the theoretical optimal.
-The higher peak demand and energy consumption presented greater challenges for optimal operations of the DER system.
-TABLE III: Summary of different control strategies under test cases
+Furthermore, for almost all cases, control algorithms showed better performance for the smaller building of HouseZero, with LSPS performing only 4.52% below the theoretical optimal.
+The performance gap between the HouseZero and BDGP buildings is mainly attributed to differences in their peak demand, total energy consumption, and net load profiles.
+Specifically, the shape of the net load depends on whether the building transitions between net-producing and net-consuming regimes, the timing of these transitions, and the duration of each regime.
+In the case of HouseZero, lower demand and higher on-site generation result in negative net load during daytime hours, with peak demand occurring in the early morning or late evening.
+In contrast, the substantially higher demand of the BDGP building prevents it from reaching a net-producing state, leading to peak demand periods during daytime hours.
+TABLE III: Summary of solution gap (%) of different control strategies
 Case Building Percentage gaps to optimal
-Backup Threshold RL LSPS
+Backup RATP RL LSPS
 Battery capacity HouseZero 47.44% 14.17% 9.33% 1.03%
@@
 Illustrated in Fig.
-3, for the first group of test cases varying battery capacity sees the LSPS algorithm performed closest to the theoretical upper bound for both HouseZero and BDGP in Fig.
+3, for the first group of test cases varying battery capacity, the LSPS algorithm performed closest to the theoretical upper bound for both HouseZero and BDGP in Fig.
 3(a) and Fig.
@@
 For the HouseZero case with low loads and large battery capacities, the peak was completely removed (0 kW), leading the LSPS algorithm to perform as well as the theoretical optimal.
-Large battery capacity also significantly improved the threshold algorithm performance, even though it significantly fell behind RL and LSPS algorithms when the battery size was more limited.
+Large battery capacity also significantly improved the RATP algorithm performance, even though it significantly fell behind the RL and LSPS algorithms when the battery size was more limited.
 On the other hand, RL followed a similar trend as the LSPS algorithm while performing around 10% worse over all cases.
-RL was still stable and more adaptive to changes in the environment than the threshold and backup algorithms.
-For the much larger BDGP building, both RL and LSPS algorithms improved overall with the increased battery capacity while backup and threshold failed to show the expected improvement due to their inflexibility in considering the high peak and energy use in this building.
+RL was still stable and more adaptive to changes in the environment than the RATP and backup algorithms.
+For the much larger BDGP building, both RL and LSPS algorithms improved overall with the increased battery capacity, while backup and RATP failed to show the expected improvement due to their inflexibility in considering the high peak and energy use in this building.
 Second, varying salvage value rate again showed the superior performance of the LSPS algorithm for both HouseZero and BDGP in Fig.
@@
 3(f).
-The threshold algorithm tended to use up the available battery charge, leading to low salvage value at the end of the day and causing a larger gap to optimal with increased salvage value rate.
+The RATP algorithm tended to use up the available battery charge, leading to low salvage value at the end of the day and causing a larger gap to optimal with an increased salvage value rate.
 Meanwhile, for extremely high salvage value cases, the backup case got closer to optimal as the strategy kept the battery full at the end of the day.
 RL performed relatively well and was closer to the LSPS algorithm and optimal for the relatively smaller HouseZero building.
-Overall, for both HouseZero and BDGP, the performances exhibited a similar trend across the four control strategies/algorithms.
-Third, varying export rates highlighted the difference between HouseZero and BDGP buildings illustrated in Fig.
+Overall, for both HouseZero and BDGP, the performances exhibited a similar trend across the four control algorithms.
+Third, varying export rates highlighted the difference between HouseZero and BDGP buildings, illustrated in Fig.
 3(c) and Fig.
 3(g).
-For the BDGP building, the higher energy demand meant the building would never be in net producing regime, leading to flat lines for the control strategies/algorithms.
-With RL, there were slight variations in performance from the stochasticity during training.
-For HouseZero, all control strategies/algorithms improved with a higher export rate since it generated more revenue for the building to sell back excessive energy production from solar.
+For the BDGP building, the higher energy demand meant the building would never be in a net producing regime, leading to flat lines for the control algorithms.
+With RL, there were slight variations in performance due to the stochasticity during training.
+For HouseZero, all control algorithms improved with a higher export rate since it generated more revenue for the building to sell back the excess energy production from solar.
 LSPS and RL algorithms performed relatively well and were stable across different export rates over the two test buildings.
-Lastly, varying peak demand charges saw all control strategies/algorithms change in a similar trend shown in Fig.
+Lastly, varying peak demand charges saw all control algorithms change in a similar trend shown in Fig.
 3(d) and Fig.
 3(h).
-It was observed that for low peak charges, all control strategies/algorithms performed well, highlighting the complexity of peak demand charges added to optimally managing such a DER system.
-With higher peak demand charges, all control strategies/algorithms diverged from optimal, with LSPS and RL algorithms performing the best across different scenarios.
+It was observed that for low peak charges, all control algorithms performed well, highlighting the complexity of peak demand charges added to optimally managing such a DER system.
+With higher peak demand charges, all control algorithms diverged from optimal, with LSPS and RL algorithms performing the best across different scenarios.
 Additionally, both of them showed a smaller solution gap for the smaller building of HouseZero, indicating the challenges for buildings with higher demand and energy consumption.
+In addition to analyzing the sensitivity of the above-mentioned parameters, we compared the performance of the RL and LSPS algorithms under different load and renewable patterns in Tab.
+IV.
+Results show that higher generation and lower demand provided higher surplus and a smaller gap to optimal from both RL and LSPS.
+On average, RL performed with a gap of 17.93% while LSPS performed with a gap of 4.43%.
+The results demonstrated the effectiveness of the RL and LSPS across various load patterns, highlighting the potential for implementing the algorithm in future long-term real-world tests.
+TABLE IV: RL and LSPS performance under different generation and demand scenarios.
+Surplus Gap to Optimal
+Generation Demand RL LSPS Optimal RL LSPS
+75th Percentile 25th Percentile 16.4 18.9 20.3 19.2% 6.9%
+50th Percentile 50th Percentile 18.2 21.0 21.9 16.9% 4.1%
+25th Percentile 75th Percentile 18.1 21.5 22.0 17.7% 2.3%
 Overall, both LSPS and RL showed great robustness and performance across different scenarios with varying battery capacities, salvage values, and prices of electricity.
-Specifically, LSPS optimization outperformed backup and threshold algorithms by 68.47% and 68.9% while RL outperformed them by 52.49% and 48.01% respectively.
-LSPS and RL algorithms trended in sync in their gap to optimality, demonstrating their ability to capture system dynamics while the LSPS algorithm is superior in around 81% instances.
-Moreover, LSPS was optimal with a large battery capacity and close to optimal for low export rate and peak demand charge.
-With 0 or low peak demand charges, all control strategies/algorithms performed close to optimal.
-In comparison, the threshold algorithm and backup mode were sensitive to changes in battery capacity, salvage rate, peak demand charge, and export rate when the building entered the net-producing regime.
-For cases where the battery capacity was large or the peak price was low, the threshold algorithm outperformed RL and approaches optimal with deterministic predictions.
-However, with a high salvage value and demand price, the threshold algorithm’s performance degraded significantly and became even worse than the backup mode.
+Specifically, LSPS optimization outperformed backup and RATP algorithms by 68.47% and 68.9%, while RL outperformed them by 52.49% and 48.01%, respectively.
+LSPS and RL algorithms trended in sync in their gap to optimality, demonstrating their ability to capture system dynamics, while the LSPS algorithm is superior in around 81% instances.
+Moreover, LSPS was optimal with a large battery capacity and close to optimal for a low export rate and peak demand charge.
+With 0 or low peak demand charges, all control algorithms performed close to optimal.
+In comparison, the RATP and backup algorithms were sensitive to changes in battery capacity, salvage rate, peak demand charge, and export rate when the building entered the net-producing regime.
+For cases where the battery capacity was large or the peak price was low, RATP outperformed RL and approached optimal with deterministic predictions.
+However, with a high salvage value and demand price, the RATP algorithm’s performance degraded significantly and became even worse than the backup mode.
 Backup mode varied mostly linearly in its gap to the optimal case.
@@
 As an example, Fig.
-4 showed the action sequence for the baseline configuration for HouseZero over one test day.
-All strategies followed a similar behavior which discharged the battery at the beginning and end of the day.
-At the same time, demand was reduced over the evening time when there were no solar generations.
-Both threshold and RL were more aggressive in battery charging actions when lacking generation predictions.
-The optimal battery actions indicated a much slower charging behavior over the daytime with precise knowledge of solar radiation, creating the performance gap mentioned above.
+4 shows the action sequence for the baseline configuration of HouseZero over one test day.
+All strategies followed a similar behavior, which discharged the battery at the beginning and end of the day.
+At the same time, demand was reduced in the evening when there was no solar generation.
+Both RATP and RL were more aggressive in battery charging actions when lacking generation predictions.
+The optimal battery actions indicated a much slower charging behavior over the daytime with precise knowledge of solar generation, creating the performance gap mentioned above.
 It was also observed that the LSPS algorithm may generate aggressive demand reduction actions due to the lack of predictions.
-Such behavior negatively impacted the utility function and the overall objective, causing gaps in the optimal solutions.
+Such behavior negatively impacted the utility function and the overall objective, causing larger solution gaps.
 VI Conclusion
@@
 To address computational complexity, we develop an efficient demand peak searching approximation method, based on a relaxed version of the problem, leading to a linear-complexity algorithm that identifies the optimal peak demand.
-Through simulations with real data from two rich datasets, we demonstrated the performance of the LSPS algorithm in achieving a small solution gap compared to baseline approaches, highlighting its effectiveness compared to relatively more computationally prohibitive methods such as model-predictive control and RL.
+Through simulations with real data from two rich and distinct datasets, we demonstrated the performance of the LSPS algorithm in achieving a small solution gap compared to baseline approaches, highlighting its effectiveness compared to relatively more computationally prohibitive methods such as model-predictive control and RL.
 The work presented here has several limitations that open avenues for future research.
@@
 Designing a more sophisticated algorithm that reliably converges to the theoretical upper bound remains an important direction for future work.
-Second, the proposed algorithm assumes full knowledge of the utility’s functional form $U(\cdot)$, which may be difficult to obtain in practice.
+Second, the LSPS algorithm assumes full knowledge of the utility’s functional form $U(\cdot)$, which may be difficult to obtain in practice.
 This limitation can be addressed using machine learning or RL techniques that infer the utility function from observed behavior [35].
+Lastly, it is worthwhile to investigate whether sufficient conditions can be established under which the LSPS algorithm remains optimal, even when projections onto the feasible solution set are involved.